### v1

# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

Abstract
This paper proposes a novel approach for statistical modelling of a continuous random variable $X$ on $[0,1)$, based on its digit representation $X=.X_{1}X_{2}\ldots$. In general, $X$ can be coupled with a random variable $N$ so that if a prior of $N$ is imposed, $(X_{1},\ldots,X_{N})$ becomes a sufficient statistics and .$X_{N+1}X_{N+2}\ldots$ is uniformly distributed. In line with this fact, and focusing on binary digits for simplicity, we propose a family of generalized finite Pólya trees that induces a random density for a sample, which becomes a flexible tool for density estimation. Here, the digit system may be random and learned from the data. We provide a detailed Bayesian analysis, including closed form expression for the posterior distribution which sidesteps the need of MCMC methods for posterior inference. We analyse the frequentist properties as the sample size increases, and provide sufficient conditions for consistency of the posterior distributions of the random density and $N$. We consider an extension to data spanning multiple orders of magnitude, and propose a prior distribution that encodes the so-called extended Newcomb-Benford law. Such a model shows promising results for density estimation of human-activity data. Our methodology is illustrated on several synthetic and real datasets.
Keywords: binary and decimal numeral systems, coupling, density estimation, extended Newcomb-Benford law, frequentist properties of Bayesian estimators, random nested partitions, round-off error.
1 Introduction
Consider a continuous random variable $X$ defined on the unit interval and let $X_{1},X_{2},\ldots$ be its binary digits such that $X=\sum_{n=1}^{\infty}X_{n}2^{-n}$. In practice, due to a round-off error or the finite arithmetic precision of computers, the data consist only of a finite number of digits. In fact, under very weak conditions, there is a coupling of $X$ with a non-negative integer-valued random variable $N$ such that $X_{1},\ldots,X_{N}$ become the sufficient digits in the sense of Møller [29] and they are independent of the remaining digits $X_{N+1},X_{N+2},\ldots$ which contain no information.
1.1 Overview on models and results
In this paper, we build on this coupling result from [29] and propose a novel approach to statistical modelling for a random variable based on its digit representation. As detailed in Theorem 3.2 and Remark 4, the coupling result can be presented in a very general setting with random digits defined in terms of nested partitions of the unit interval, where the intervals of each partition are not necessarily of equal length as explained in Definition 3.1 and Remark 2 below. Thinking of $N$ as an latent variable which determines the needed number of digits, and working with binary digits for simplicity, conditioned on $N=n$, we propose models where for $j=1,\ldots,n$, we let $X_{j}\,|\,X_{1},\ldots,X_{j-1}$ be Bernoulli-distributed with a beta-distributed probability parameter depending on $X_{1},\ldots,X_{j-1}$, and where $X_{n+1},X_{n+2},\ldots$ are independent of the first $n$ digits and contain no information. Our prior models can be seen as generalizations of Pólya trees (PT) [25, 26, 27] with a finite number of partition levels, and are therefore termed generalized finite Pólya trees of type 1 (for short GFPT1; we will introduce the type 2 models below). PT priors play an important role in non-parametric Bayesian inference, see e.g. [1, 6, 7, 16, 18, 23, 31, 33, 38]. A GFPT1 prior specifies the distribution of a random probability density function (PDF) used for the data model of exchangeable observations. This random PDF is a mixture over $N$ of functions on the intervals of the binary partitions.
We provide a closed-form expression for the posterior distribution which allows for straightforward inference without resorting to Markov chain Monte Carlo algorithms. In fact, our models can be fitted to tens of millions of datapoints in a matter of a fraction of a second on a standard laptop. We investigate the frequentist properties of the Bayesian estimators obtained from the posterior distribution when assuming each data point was generated from a true PDF $f^{*}$: First, we consider the posterior distribution of $N$. Assuming that $n^{*}$ digits are sufficient for $f^{*}$ (in the sense of Theorem 3.2), we establish conditions under which the posterior of $N$ concentrates on $n^{*}$, assuming that we have observed either the data perfectly (i.e., without round-off errors), or that we have access only to the first $\bar{n}$ digits, which is arguably the most common situation in applied settings. In the latter case, we identify a curious situation where ignoring the round-off errors leads to $N$ diverging to infinity a posteriori. Instead, assuming that the round-off error is uniform leads to the correct asymptotic behaviour. Second, we consider the convergence of the random PDF, establishing consistency with respect to both the Kolmogorov-Smirnov and the Hellinger distance.
We propose two generalizations of GFPT1 priors. For the first ones, termed generalized finite Pólya trees of type 2 (GFPT2), we let the nested binary partitions be random such that the lengths of the intervals are obtained from sequences of beta-distributed random variables as detailed later in Definition 4.5. This additional stochastic layer is essential: Under a fixed deterministic partition one may require an unbounded number of digits ($n^{\ast}=\infty$) even when the true PDF $f^{\ast}$ is merely piecewise constant, simply because its discontinuities do not coincide with the pre-specified nested partitions. By randomising the partitions, a GFPT2 prior circumvents this limitation and restores the finite digit sufficiency property for a broad class of densities. The second generalization, discussed in Section 6, proposes how to model a sample consisting of positive numbers by combining a GFPT1 prior with a prior distribution for the order of magnitude of each data point. We find such a model to be particularly useful for density estimation in problems where data span multiple orders of magnitude and obey the so-called extended Newcomb-Benford law which routinely happens in human data [5, 8, 22]. In brief, the Newcomb-Benford law asserts that small digits are more likely to occur as leading digit in human data, and we show in Section 4.3 how to incorporate this as prior information by extending certain GFPT1s to what we call multiscale Benford Pólya trees (MBPT).
We mainly restrict attention to the case $10<X<$ 1, since any real number can be transformed to a number between 0 and 1 by a bijective mapping (e.g. $x\mapsto\exp(x)/(1+\exp(x))$) and since data consisting of positive numbers can be modelled by using MBPT priors. To cover data consisting of real numbers, MBPT priors may of course be extended by including a prior for the sign.
1.2 General numeral systems
It is only for specificity and simplicity that we consider a binary numeral system. Indeed, everything in this paper easily extends to the case of base-$q$ representations of numbers when $q\geq 2$ is an integer (further details are given in Section 4.3). This includes the common cases of a binary ($q=2$) and a decimal numeral system ($q=10$) with equal interval lengths in each partition, and for applications it seems not important how $q$ is specified: Suppose $N_{q}$ is the discrete random variable corresponding to $N_{2}=N$ above but obtained by using base-$q$ digits instead, whereby $q^{N_{q}}$ becomes the number of possible values of the sufficient digits $X_{1},\ldots,X_{N_{q}}$ when the value of $N_{q}$ is fixed. Considering simulation studies and statistical analysis of real data examples, applying our prior models such that $2^{N_{2}}$ and $10^{N_{10}}$ are closely distributed, we observe in Section 7.3 that rather similar conclusions are obtained no matter if $q=2$ or $q=10$.
1.3 Outline
The paper is organized as follows. Section 2 specifies the notation used throughout this text. Section 3 explains the meaning of sufficient digits (as briefly discussed in Section 1.1). Models for data on the unit interval are studied in Section 4: Section 4.1 recalls the definition of Pólya tree priors and their finite versions, Section 4.2 introduces our generalized finite Pólya tree priors GFPT1 and GFPT2 (cf. Section 1.1), and Section 4.3 discusses the choice of hyperparameters for such priors and the common Pólya tree priors. Various theoretical results for Bayesian analysis based on GFPT1 and GFPT2 priors are established in Section 5: Section 5.1 shows how to easily simulate from the posterior distribution. Mainly restricting attention to GFPT1 priors and shortly commenting on GFPT2 priors, Section 5.2 deals with consistency results for the posterior distribution of $N$ and Section 5.3 with consistency results for the posterior density estimate (as briefly discussed in Section 1.1). Section 6 studies our generalization of GFPT1 priors via the extended Newcomb-Benford law to MBPT priors, cf. Section 1.1. Section 7 considers numerical simulation results: Section 7.1 shows how the choice of hyperparameters affects posterior inference when using GFPT1 priors and suggests reasonable default values. Section 7.2 compares the posterior density estimate when using PT and GFPT1 priors and assess whether the increased flexibility of GFPT2 priors yields practical gains over GFPT1 ones. Section 7.3 compares results for Bayesian density estimation when using GFPT1 and GFPT2 priors, base-2 and base-10 MBPT priors, and the standard prior given by a Dirichlet process mixture of Gaussian densities (see e.g. [14]). Finally, Section 8 summarizes our findings and points to future research.
The Supplemental material collects the missing proofs in the present paper.
Supporting Julia code implementing the proposed methodology and the numerical simulations is available at https://github.com/mberaha/FinitePolyaTrees.git
2 Notation
Let =$\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{N}_{0}=\{0,1,2,\ldots\}$, $\mathbb{R}=(-\infty,\infty)$, and $\mathbb{R}_{+}=(0,+\infty)$ be the sets of positive integers, non-negative integers, real numbers, and positive real numbers, respectively.
For $n\in\mathbb{N}$ and $(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, let $x_{1:n}=(x_{1},\ldots,x_{n})$, and for $k\in\{0,1\}$, identify $x_{1:n-1},k$ by $(x_{1},\ldots,x_{n-1},k)$ if $n>1$ and by $k$ if $n=1$. Set $x_{1:0}=\emptyset$, $\{0,1\}^{0}=\{\emptyset\}$, and $x_{\emptyset,k}=x_{k}$ for $k=0,1$. We may interpret $x_{1:n}$ as an ordered configuration of $n$ points if $n>0$ and as the empty point configuration if $n=0$ (this is similar to the terminology used in point process theory [30]).
The lexicographic order for finite binary sequences of possibly different lengths, denoted by $\prec$, is defined as follows. For $m,n\in\mathbb{N}$, $x_{1:n}\in\{0,1\}^{n}$, and $z_{1:m}\in\{0,1\}^{m}$, write $x_{1:n}\prec z_{1:m}$ if either $m\geq n$ and $z_{1:n}=x_{1:n}$, or there exists a non-negative integer $k<\min\{m,n\}$ such that $x_{1:k}=z_{1:k}$, $x_{k+1}=0$, and $z_{k+1}=1$.
Denote $\{0,1\}^{\mathbb{N}}$ the set of all infinite sequences $(x_{1},x_{2},\ldots)$ with $x_{n}\in\{0,1\},n=1,2,\ldots$. Let $\mathcal{I}$ be the collection of all half-open intervals $[a,b)$ with $10\leq a<b<$ 1 (where $a$ is included and $b$ is excluded). For a set $A$ with either $A=\mathcal{I}$ or $A\subseteq\mathbb{R}$, for $(x_{1},x_{2},\ldots)\in\{0,1\}^{\mathbb{N}}$, and for intervals/numbers $z_{x_{1:j}}\in A$ with $j\in\mathbb{N}$, consider $z=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},\,j\in\mathbb{N})$ as an infinite sequence with entries $z_{x_{1:j}}$ which appear in accordance to the ordering of the $x_{1:j}$ with respect to $\prec$. For $n\in\mathbb{N}$, restricting $z$ to its first $2^{n}$ entries, we obtain the ordered finite sequence $z^{(n)}=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},\,j=1,\ldots,n)$. For notional convenience, let $z^{(n)}=\emptyset$ denote the empty set (sequence). For $n\in\mathbb{N}_{0}$, let $z^{(>n)}=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},j=n+1,n+2,\ldots)$. Thus, $z^{(>0)}=z$.
Denote the beta-distribution on $[0,1)$ with shape parameters $\alpha_{0}>0$ and $\alpha_{1}>0$ by $\mathrm{B}(\alpha_{0},\alpha_{1})$, and its PDF by $\mathrm{B}(\cdot\,|\,\alpha_{0},\alpha_{1})$. For later use, extend this PDF to $[0,1]$ by setting $\mathrm{B}(1\,|\,\alpha_{0},\alpha_{1})=0$ (or any other arbitrary value). For $k\in\{0,1\}$, $\alpha_{k}=0$, and $\alpha_{1-k}>0$, let $\mathrm{B}(\alpha_{0},\alpha_{1})$ be the degenerated probability distribution on $[0,1]$ which is concentrated at the point $k$ (so the value of $\alpha_{1-k}$ plays no role), and for $x\in[0,1]\setminus\{k\}$, let $\mathrm{B}(k\,|\,\alpha_{0},\alpha_{1})=1$ and $\mathrm{B}(x\,|\,\alpha_{0},\alpha_{1})=0$.
For numbers $x>0$ and $y>0$, the beta function is given by $\mathrm{b}(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}\,\mathrm{d}t$, and we let $b(0,x)=b(x,0)=1$.
For numbers $a<b$, $\mathrm{Unif}[a,b)$ denotes the uniform distribution on $[a,b)$.
Finally, set $10^{0}=$ 1.
3 The sufficient digits
The material in this section is used to motivate our new models introduced in Section 4. For the following definition, let $a_{\emptyset}=0$ and length $\ell_{\emptyset}=1$.
Definition 3.1.
Suppose $I=(I_{x_{1:n}}\,|\,x_{1:n}\in\{0,1\}^{n},\,n\in\mathbb{N})$ where each $I_{x_{1:n}}=[a_{x_{1:n}},a_{x_{1:n}}+\ell_{x_{1:n}})$ has left endpoint $a_{x_{1:n}}\in[0,1)$ and length $\ell_{x_{1:n}}\in(0,1)$ such that for every $n\in\mathbb{N}_{0}$ and every $x_{1:n}\in\{0,1\}^{n}$,
$a_{x_{1:n}}=a_{x_{1:n},0},\quad a_{x_{1:n},1}=a_{x_{1:n},0}+\ell_{x_{1:n},0},\quad\ell_{x_{1:n}}=\ell_{x_{1:n},0}+\ell_{x_{1:n},1}.$ (1)
We call $I$ a NBP of $[0,1)$ (this abbreviation is explained in Remark 1 below). Finally, define the digits of every $x\in[0,1)$ as the unique sequence $(x_{1},x_{2},\ldots)\in\{0,1\}^{\mathbb{N}}$ such that $x\in I_{x_{1:n}}$ for $n=1,2,\ldots$.
For instance, for the usual binary numeral system, $a_{x_{1:n}}=\sum_{i=1}^{n}x_{i}2^{-i}$ and $\ell_{x_{1:n}}=2^{-n}$. Then we refer to $I$ as the standard diadic partitions.
Remark 1.
Let the situation be as in Definition 3.1. For all $x_{1:n}\in\{0,1\}^{n}$ with $n\in\mathbb{N}_{0}$, the intervals $I_{x_{1:n},0}$ and $I_{x_{1:n},1}$ are disjoint, $a_{x_{1:n},1}$ is the right endpoint of $I_{x_{1:n},0}$ and the left endpoint of $I_{x_{1:n},1}$, and $I_{x_{1:n}}=I_{x_{1:n},0}\cup I_{x_{1:n},1}$. We call $I_{x_{1:n},0}$ the left interval of this split. For every $n\in\mathbb{N}$,
$[0,1)=\bigcup_{x_{1:n}\in\{0,1\}^{n}}I_{x_{1:n}},$ (2)
where the $2^{n}$ sets on the right hand side are pairwise disjoint sets, and so by (1), the infinite sequence $I^{(1)},I^{(2)},\ldots$ constitutes a collection of nested binary partitions of $[0,1)$ (explaining the abbreviation NBP). By (1), the left endpoints of the intervals in $I$ are ordered in accordance to $\prec$, i.e., for any $x_{1:n}\in\{0,1\}^{n}$ and $z_{1:m}\in\{0,1\}^{m}$ with $n,m\in\mathbb{N}$ and $x_{1:n}\prec z_{1:m}$, we have $a_{x_{1:n}}<a_{z_{1:m}}$. By (1), for every $n\in\mathbb{N}$, $I^{(n)}$ is determined by $(\ell_{x_{1:j-1},0}\,|\,x_{1:j}\in\{0,1\}^{j},j=1,\ldots,n)$ which is the ordered sequence of the lengths of the left intervals up to level $n$.
Remark 2.
Henceforth, assume for every $x\in[0,1)$ with digits $(x_{1},x_{2},\ldots)$ that $\ell_{x_{1:n}}\to 0$ as $n\to\infty$. Hence, $x$ is determined by its digits and there is a one-to-one correspondence between $[0,1)$ and $\{0,1\}^{\mathbb{N}}$. We express this by writing $x=.x_{1}x_{2}\ldots$.
Definition 3.1 may be extended by allowing an interval $I_{x_{1:n}}$ to be empty. Such an extension is relevant for certain number systems, e.g. in relation to so-called pseudo $\beta$-expansions, cf. [21] and the references therein. It is rather straightforward to modify the ideas and results in this paper to such situations.
Remark 3.
The remainder of this paper considers the following setting. Let $X$ be a random variable with a PDF $f$ concentrated on $H\subseteq[0,1)$ where $H$ has Lebesgue measure 1, and let $X_{1},X_{2},\ldots$ be the random digits of $X=.X_{1}X_{2}\ldots$. Assume that $f$ is lower semi-continuous (LSC) on $H$. Indeed, this is a very mild condition, cf. [20].
We need some notation for the following theorem. Define $c_{\emptyset}=\inf_{H}f$. For $n\in\mathbb{N}$ and $x_{1:n}\in\{0,1\}^{n}$, let $i_{{1:n}}=\inf_{H\cap I_{x_{1:n}}}f$ and $c_{x_{1:n}}=i_{{1:n}}-i_{{1:n-1}}$, where we let the infimum over the empty set be 0. Denote $2^{\mathbb{N}_{0}}$ the set of all subsets of $\mathbb{N}_{0}$, and $\mathcal{B}$ the set of Borel sets included in $H$. Equip the product space $H\times\mathbb{N}_{0}$ with the product $\sigma$-algebra of $\mathcal{B}$ and $2^{\mathbb{N}_{0}}$. Let $\eta$ be the product measure of Lebesgue measure on $\mathcal{B}$ and counting measure on $2^{\mathbb{N}_{0}}$. We use the abbreviation PMF for a probability mass function.
Theorem 3.2.
Under the conditions above, there is a coupling of $X$ with a random variable $N\in\mathbb{N}_{0}$ such that the distribution of $(X,N)$ is absolutely continuous with respect to $\eta$, with a density for any $(x,n)\in H\times\mathbb{N}_{0}$ given by
$f(x,n)=c_{x_{1:n}}\quad\mbox{if }x=.x_{1}x_{2}\ldots$ (3)
Conversely, suppose $Q$ is a probability measure on $H\times\mathbb{N}_{0}$ which is absolutely continuous with respect to $\eta$ such that its density $q={\mathrm{d}Q}/{\mathrm{d}\eta}$ is of the form
$q(x,n)=g_{n}(x)\quad\mbox{for all }(x,n)\in H\times\mathbb{N}_{0}$
where each $g_{n}$ is a non-negative LSC function on $H$ with finite Lebesgue measure. Defining the PDF
$g(x)=\sum_{n=0}^{\infty}g_{n}(x)\quad\mbox{for all }x\in H$
and letting $N$ be a random variable with PMF $p_{n}=\int_{H}g_{n}$ for $n\in\mathbb{N}_{0}$, then $g$ is LSC on $H$ and $X\,|\,N=n$ has PDF $f_{n}=g_{n}/p_{n}$ for $p_{n}>0$. In particular, (3) is the special case where each $g_{n}$ is a non-negative constant function on each interval $I_{x_{1:n}}$ such that $\sum_{n=0}^{\infty}\int_{0}^{1}g_{n}(x)\,\mathrm{d}x=1$.
The proof of Theorem 3.2 is found in the Supplemental material.
Remark 4.
Clearly, $(X,N)$ and $(X_{1:N},.X_{N+1}X_{N+2}\ldots)$ are in a one-to-one correspondence. Thinking of $X$ as data and $N$ as a latent variable, considering a statistical model of $f$ and imaging we could observe the missing data $N$, (3) shows that $X_{1:N}$ is a sufficient statistic with distribution
$\mathrm{P}(X_{1:N}=x_{1:n})=\ell_{x_{1:n}}c_{x_{1:n}}\quad\mbox{for every $n\in\mathbb{N}_{0}$ and every $x_{1:n}\in\{0,1\}^{n}$.}$
On the other hand, the remainder .$X_{N+1}X_{N+2}\ldots$ is an ancillary statistic, and it is shown in [29] (see also [19, 20]) that
.$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:N}$. (4)
The second part of Theorem 3.2 specifies a more general structure, where it would be interesting to see how the ideas introduced in the present paper might be extended.
We interpret $X_{1:N}$ as a finite ordered point process, where $N$ is the random number of points and the state space is $\Omega=\cup_{n=0}^{\infty}\{0,1\}^{n}$. Equip $\Omega$ with the $\sigma$-algebra $\mathcal{G}$ generated by the sets $A\subseteq\{0,1\}^{n}$ with $n\in\mathbb{N}_{0}$. Let $\mathcal{P}_{\mathrm{FOPP}}$ denote the class of all probability distributions on $(\Omega,\mathcal{G})$ and $\mathcal{P}_{\mathrm{LSC}}$ the class of all absolutely continuous probability distributions on $[0,1)$ with a density which is almost everywhere LSC. Theorem 3.2 gives immediately the following corollaries.
Corollary 3.3.
The coupling construction of $(X,N)$ in Theorem 3.2 establishes a one-to-one correspondence between $\mathcal{P}_{\mathrm{FOPP}}$ and $\mathcal{P}_{\mathrm{LSC}}$.
Remark 5.
Henceforth, since $N$ is not observable, we impose a prior PMF
$p_{n}=\mathrm{P}(N=n),\quad n\in\mathbb{N}_{0}.$
Corollary 3.4.
Under the coupling construction of $(X,N)$ in Theorem 3.2, $X$ conditioned on $N$ has PDF
$f(x\,|\,N=n)=\mathrm{P}(X_{1:n}=x_{1:n})/{\ell_{x_{1:n}}}\quad\mbox{if $x=.x_{1}x_{2}\ldots\in[0,1)$ and $n\in\mathbb{N}_{0}$.}$
Corollary 3.4 is in accordance with (4) and to complete the description of the distribution of $X$ it remains only for every $n\in\mathbb{N}$ to specify the PMF $\mathrm{P}(X_{1:n}=x_{1:n})$ for $x_{1:n}\in\{0,1\}^{n}$. This is the subject of Section 4.
4 Models
4.1 Finite Pólya tree distributions
Definition 4.1.
Let $\alpha=(\alpha_{x_{1:n}}\,|\,x_{1:n}\in\{0,1\}^{n},\,n\in\mathbb{N})$ be a given parameter such that $\alpha_{x_{1:n}}\geq 0$ and $\alpha_{x_{1:n-1},0}+\alpha_{x_{1:n-1},1}>0$ whenever $x_{1:n}\in\{0,1\}^{n}$ and $n\in\mathbb{N}$. Suppose
$Y=(Y_{x_{1:n}}\,|\,x_{1:n}\in\{0,1\}^{n},\,n\in\mathbb{N})$
is a stochastic process with each $Y_{x_{1:n-1},0}\sim\mathrm{B}(\alpha_{x_{1:n-1},0},\alpha_{x_{1:j-1},1})$ and $Y_{x_{1:j-1},1}=1-Y_{x_{1:j-1},0}$, and where the $Y_{x_{1:n-1},0}$’s with $n\in\mathbb{N}$ are independent. Then $Y$ is said to be an infinite beta-distributed sequence with parameter $\alpha$, and for any $n\in\mathbb{N}$, $Y^{(n)}$ to be a finite beta-distributed sequence with parameter $\alpha^{(n)}$ and $Y^{(>n)}$ to be an infinite beta-distributed sequence with parameter $\alpha^{(>n)}$. For short, write
$Y\sim\mathrm{BS}(\alpha),\quad Y^{(n)}\sim\mathrm{FBS}(\alpha^{(n)}),\quad Y^{(>n)}\sim\mathrm{BS}(\alpha^{(>n)}).$
The finite beta-distributed sequence induces a random PDF
$f\big{(}x\,|\,Y^{(n)}\big{)}=\frac{\prod_{j=1}^{n}Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}}{\ell_{x_{1:n}}}\quad\mbox{if $x=.x_{1}x_{2}\ldots\in[0,1)$}$ (5)
and we write
$f\big{(}\cdot\,|\,Y^{(n)}\big{)}\sim\Pi^{(n)}=\mathrm{FPT}(\alpha^{(n)},I^{(n)})$
for its distribution which is called a finite Pólya tree with parameter $(\alpha^{(n)},I^{(n)})$.
Remark 6.
The random PDF $f\big{(}\cdot\,|\,Y^{(n)}\big{)}$ is constant on every interval $I_{x_{1:n}}$ with $x_{1:n}\in\{0,1\}^{n}$. Using $\Pi^{(n)}$ as a prior distribution generates a distribution of $X=.X_{1}X_{2}\ldots$ where $X\,|\,Y^{(n)}\sim f(\dot{\,}|\,Y^{(n)})$. Then by (5), for $j=1,\ldots,n$, conditioned on both $Y^{(n)}$ and $X_{1:j-1}$, the probability that $X_{j}=0$ is $Y_{X_{1:j-1},0}$, and .$X_{n+1}X_{n+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:n}$. Recall that $\alpha_{x_{1:j-1},x_{j}}$ is allowed to be 0: If $\alpha_{x_{1:j-1},x_{j}}=0$ then $Y_{x_{1:j-1},x_{j}}=x_{j}$ (see Section 2) which implies $X_{j}=1-x_{j}$
Definition 4.2.
Letting $n\to\infty$ in Definition 4.1 , we recover the definition of a Pólya tree $\Pi^{(\infty)}=\mathrm{PT}(\alpha,I)$ as given by Lavine [25]. We refer to $\Pi^{(\infty)}$ as the PT prior.
Remark 7.
Arguments for considering a finite Pólya tree distribution can be found in [26, 38]. For instance, $\Pi^{(\infty)}$ is not always the distribution of a random PDF unless the $\alpha$’s increase sufficiently rapidly which leads to problems of robustness, cf. [26]. Further, if $\Pi^{(\infty)}$ is the distribution of a random PDF, this random PDF is almost surely discontinuous almost everywhere, cf. [12].
Definition 4.3.
Let $\mathrm{FBS}(\alpha^{(0)})$ denote the degenerated distribution concentrated at $Y^{(0)}=\emptyset$. In accordance with Corollary 3.4, extend (5) to the case $n=0$ by setting $\prod_{j=1}^{n}\cdots=1$ so that $f\big{(}x\,|\,Y^{(0)}\big{)}=1$ is the uniform PDF and $\Pi^{(0)}=\mathrm{FPT}(\alpha^{(0)},I^{(0)})=\{f\big{(}\cdot\,|\,Y^{(0)}\big{)}\}$. Here, $\alpha^{(0)}$ and $I^{(0)}$ have no meaning and are just introduced for notional convenience.
4.2 Generalized finite Pólya tree distributions
This section generalizes the finite Pólya tree distribution in two directions by assuming that the truncation level $n$ is random and possibly also the nested binary partitions $I^{(1)},I^{(2)},\ldots$ are random. The motivation for using a random truncation level $N$ is given by Theorem 3.2 and Corollary 3.3. Note that $n\in\mathbb{N}_{0}$ is determined by $Y^{(n)}$, since $Y^{(n)}$ has dimension $2^{n}$.
Definition 4.4.
Given a NBP $I$ as in Definition 3.1, an infinite beta-distributed sequence $Y\sim\mathrm{BS}(\alpha)$ as in Definition 4.1, $Y_{0}=\emptyset$ as in Definition 4.3, and a discrete random variable $N$ on $\mathbb{N}_{0}$ which is independent of $Y$, then $(N,Y)$ induces a random PDF depending only on $Y^{(N)}$ and given by
$f\big{(}x\,|\,Y^{(N)}\big{)}=\frac{\prod_{j=1}^{N}Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}}{\ell_{x_{1:N}}}\quad\mbox{if $x=.x_{1}x_{2}\ldots\in[0,1)$}.$ (6)
We write
$f\big{(}\cdot\,|\,Y^{(N)}\big{)}\sim\Pi_{1}=\mathrm{GFPT1}(\alpha,I)$ (7)
for its distribution which is called a generalized finite Pólya tree of type 1 with parameter $(\alpha,I)$. We refer to $f\big{(}\cdot\,|\,Y^{(N)}\big{)}$ as the random GFPT1 PDF and to $\Pi_{1}$ as the GFTP1 prior. We also refer to the distribution of $(Y,N)$ as the GFTP1 prior (it will be clear by the context what is meant).
Remark 8.
Let $X\sim f\big{(}\cdot\,|\,Y^{(N)}\big{)}\sim\Pi_{1}$. Conditioned on $(Y,N)$ (or just $Y^{(N)}$), we have that
.$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ (8)
is independent of $X_{1:N}$ which follows the random PMF given by the numerator in (6). This is in agreement with (4), and the law of $X$ can be expressed by a Bayesian hierarchical model with three steps:
$X\,|\,Y^{(N)}$ $\sim f\big{(}x\,|\,Y^{(N)}\big{)}$ (9)
$Y^{(N)}\,|\,N$ $\sim\mathrm{FBS}(\alpha^{(N)})$
$N$ $\sim p$
[…]
Given a sample $Z_{1},\ldots,Z_{m}$ which share the same latent variables $Y$ and $N$ so that
$Z_{1},\ldots,Z_{m}\,|\,Y,N\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}f\big{(}\cdot\,|\,Y^{(N)}\big{)}\sim\Pi_{1},$ (10)
the inferential goal is to derive the posterior distribution of $(Y,N)$. The posterior distribution of the random GFTP1 PDF is denoted $\Pi_{1}(\cdot\,|\,Z_{1:m})$ and called the posterior GFTP1 PDF:
$f\big{(}\cdot\,|\,Y^{(N)}\big{)}\,|\,Z_{1:m}\sim\Pi_{1}(\cdot\,|\,Z_{1:m}).$ (11)
Our objectives in this paper are the posterior GFTP1 PDF and the posterior distribution of $N$ under the GFPT1 prior.
So far the NBP $I$ has be given. To obtain further flexibility, randomness of $I$ may be imposed: For every $n\in\mathbb{N}$ and every $x_{1:n}\in\{0,1\}^{n}$, suppose $I_{x_{1:n}}$ has a random length $L_{x_{1:n}}$, let
$R_{x_{1:n}}=L_{x_{1:n}}/L_{x_{1:n-1}}$ (12)
with $L_{x_{1:0}}=1$, and let $R=(R_{x_{1:n}}\,|\,x_{1:n}\in\{0,1\}^{n},n\in\mathbb{N})$. Then there is a one-to-one correspondence between $I^{(n)}$ and $R^{(n)}$, and between $I$ and $R$. Specifically, consider the following case, where we let $\mathrm{FBS}(\beta^{(0)})$ denote the degenerated distribution concentrated at the empty set (sequence) $R^{(0)}=\emptyset$.
Definition 4.5.
Let $\beta=(\beta_{x_{1:n}}\,|\,x_{1:n}\in\{0,1\}^{n},\,n\in\mathbb{N})$ be a given parameter with each $\beta_{x_{1:n}}>0$. Suppose Definition 4.4 is extended such that $R\sim\mathrm{BS}(\beta)$ is independent of $(N,Y)$, and $(N,Y,R)$ induces a random PDF depending only on $(Y^{(N)},R^{(N)})$ and given by
$f\big{(}x\,|\,Y^{(N)},R^{(N)}\big{)}=\frac{\prod_{j=1}^{N}Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}}{L_{x_{1:N}}}\quad\mbox{if $x=.x_{1}x_{2}\ldots\in[0,1)$}$ (13)
where $x_{1:N}$ is specified by $R^{(N)}$. We write
$f\big{(}\cdot\,|\,Y^{(N)},R^{(N)}\big{)}\sim\Pi_{2}=\mathrm{GFPT2}(\alpha,\beta)$
for its distribution which is called a generalized finite Pólya tree of type 2 with parameter $(\alpha,\beta)$. We refer to $f\big{(}\cdot\,|\,Y^{(N)},R^{(N)}\big{)}$ as the random GFPT2 PDF and to $\Pi_{2}$ as the GFTP2 prior. We also refer to the distribution of $(Y,R,N)$ as the GFTP2 prior (it will be clear by the context what is meant).
Remark 9.
Let $X\sim\Pi_{2}\sim\mathrm{GFPT2}(\alpha,\beta)$. Conditioned on $(Y,R,N)$ (or just $(Y^{(N)},R^{(N)})$, .$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:N}$ which follows the random PMF given by the numerator in (13). The law of $X$ can be expressed by a Bayesian hierarchical model:
$X\,|\,Y^{(N)},R^{(N)}$ $\sim f\big{(}x\,|\,Y^{(N)},R^{(N)}\big{)}$
$Y^{(N)}\,|\,N\sim\mathrm{FBS}(\alpha^{(N)})$ $\mbox{and }R^{(N)}\,|\,N\sim\mathrm{FBS}(\beta^{(N)})\ \mbox{are independent}$
$N$ $\sim p$
[…]
Given a sample $Z_{1},\ldots,Z_{m}$ which share the same latent variables $Y$, $R$, and $N$ so that
$Z_{1},\ldots,Z_{m}\,|\,Y,R,N\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}f\big{(}\cdot\,|\,Y^{(N)},R^{(N)}\big{)}\sim\Pi_{2},$ (14)
the inferential goal is to derive the posterior distribution of $(Y,R,N)$. The posterior distribution of the GFTP2 PDF is denoted $\Pi_{2}(\cdot\,|\,Z_{1:m})$ and called the posterior GFTP2 PDF:
$f\big{(}\cdot\,|\,Y^{(N)},R^{(N)}\big{)}\,\big{|}\,Z_{1:m}\sim\Pi_{2}(\cdot\,|\,Z_{1:m}).$ (15)
In this paper, we focus on the posterior GFTP2 PDF and the posterior distribution of $N$ under the GFPT2 prior, although the posterior distribution of $R$ may possibly be of interest as well.
Remark 10.
Extending Definition 4.5 by allowing $\beta_{x_{1:n}}$ to be 0 would correspond to relaxing Definition 3.1 so that $I_{x_{1:n}}$ may be empty, in which case we should let $X_{n}=1-x_{n}$. The ideas and results in this paper may rather easily be extended to this situation.
4.3 Specification of parameters
It remains to specify the parameters $\alpha$ and $\beta$ of the PT, GFTP1, and GFTP2 priors introduced in Sections 4.1 and 4.2, and to specify a prior distribution of $N$. This section introduces some specific choices of $\alpha$ and $\beta$, which are later used in Section 7. Priors for $N$ are discussed later in Remark 16 and in Section 7.
Recalling Remark 7, following common practice, setting
$\alpha_{x_{1:n-1},0}=\alpha_{x_{1:n-1},1}=\alpha_{0}n^{2},\quad x_{1:n-1}\in\mathbb{N}^{n-1},\quad n\in\mathbb{N},$ (16)
for some $\alpha_{0}>0$, then $\Pi^{(\infty)}$ is almost surely absolutely continuous, cf. [25].
Another choice is inspired by the extended Newcomb-Benford law (the general significant-digit law of [22]): Recall that in the present paper we consider binary number representations, cf. Section 1, but the case of the usual binary numeral system extends to the general case of base-$q$ number representations with $q\geq 2$ an integer, so to make this point clear we write $q$ instead of 2 (later in Sections 6 and 7.3 we consider results for the most common cases $q=2$ and $q=10$). Consider $X=.X_{1}X_{2}\ldots=\sum_{i=1}^{\infty}X_{i}q^{-i}$, where $q^{-i}$ is the interval length of the $i$th partition of $[0,1)$. Extend $X$ by considering a continuous random variable $Z>0$ with base-$q$ number representation $Z=q^{M+1}X$, where $M$ is its order of magnitude and $X_{1}\not=0$ is its leading digit. Suppose $Z$ satisfies the extended Newcomb-Benford law (in base-$q$), which means that $\log_{q}(qX)$ is uniformly distributed between 0 and 1 – one says that $Z$ spans all orders of magnitude. Equivalently, for all $n\in\mathbb{N}$ and $x_{1:n}\in\{0,\ldots,q-1\}^{n}$,
$p_{n}(x_{1:n})=\begin{cases}\log_{q}\left[1+\left(\sum_{i=1}^{n}x_{i}q^{n-i}\right)^{-1}\right]&\text{if }x_{1}\not=0,\\ 0&\text{if }x_{1}=0.\end{cases}$ (17)
In fact many real-world datasets such as those involving incomes, city sizes, or seismic magnitudes, they span multiple orders of magnitude – see https://testingbenfordslaw.com for a list of real datasets obeying (17).
We exploit (17) to specify $\alpha$ such that the distribution of $X_{n}\,|\,X_{1:n-1}$ under the PT prior is equal to the conditional distribution of the $n$-th digit under (17). Formally, for $k=0,\ldots,q$, let
$\mathrm{P}(X_{n}=k\,|\,X_{1:n-1}=x_{1:n-1})={\mathrm{E}}[\mathrm{P}(X_{n}=k\,|\,X_{1:n-1}=x_{1:n-1},Y)]=\frac{\alpha_{x_{1:n-1},k}}{\sum_{j=0}^{q-1}\alpha_{x_{1:n-1},j}}$
and solve
$\mathrm{P}(X_{n}=k\,|\,X_{1:n-1}=x_{1:n-1})=\frac{p_{n}(x_{1:n-1},k)}{\sum_{j=1}^{q-1}p_{n}(x_{1:n-1},j)},$
which leads to
$\alpha_{x_{1:n}}=\begin{cases}c_{n}\frac{p_{n}(x_{1:n})}{\sum_{j=1}^{q-1}p_{n}(x_{1:n-1},j)}&\text{if }x_{1}\not=0,\\ 0&\text{if }x_{1}=0.\end{cases}$ (18)
Here, $(c_{n})_{n\geq 1}$ is a sequence of user-specified positive parameters that control the variance of $Y$. Specifically, $\mathrm{Var}(Y_{x_{1:n}})$ is inversely proportional to $c_{n}$. Finally, to complete the description of the distribution of $Z$, we need to specify a joint distribution of $M$ and $X$. We defer this to Sections 6 and 7.3.
Consider now a GFPT2 prior. Since we assumed every $\ell_{x_{1:n}}>0$, it is required that $\beta_{x_{1:n}}>0$ for all $n\in\mathbb{N}$ and $x_{1:n}\in\{0,1\}^{n}$. Assume ${\mathrm{E}}[L_{x_{1:n}}]=2^{-n}$, that is, the expected value of the NBP coincides with the standard diadic partitions of $[0,1)$. Hence, $\beta_{x_{1:n-1},0}=\beta_{x_{1:n-1},1}=\beta_{n}$. In our experience, values of $R_{x_{1:n}}$ too close to zero or one lead to numerical instability issues when updating $R$ via a Metropolis-Hastings algorithm as in Remark 13. Therefore, we suggest setting $\beta_{n}=2$ to avoid giving prior mass to those values of $R_{x_{1:n}}$.
5 Bayesian analysis
[…]
We use the following notation. Let the data be given by a a realization $z_{1:m}\in[0,1)^{m}$ of $Z_{1:m}$. For $j\in\mathbb{N}_{0}$ and $x_{1:j}\in\{0,1\}^{j}$, let $n_{x_{1:j}}(z_{1:m})$ be the number of observations $z_{i}$ ($i=1,\ldots,m$) falling in the interval $I_{x_{1:j}}$. By Definitions 4.4 and 4.5, we have almost surely that $n_{x_{1:j}}(Z_{1:m})=0$ if $\alpha_{x_{1:j}}=0$ and $j>1$, so assume $n_{x_{1:j}}(z_{1:m})=0$ if $\alpha_{x_{1:j}}=0$ and $j>1$.
5.1 Posterior simulation
5.1.1 Posterior simulation under a GFPT1 prior
We start by considering the posterior distribution of $(N,Y)$ for the case (10) where a GFTP1 prior has been specified. Let $\gamma(z_{1:m},\alpha)=(\gamma_{x_{1:n}}(z_{1:m},\alpha)\,|\,x_{1:j}\in\{0,1\}^{j},\,j\in\mathbb{N})$ be the infinite sequence with entries
$\gamma_{x_{1:j}}(z_{1:m},\alpha)=\alpha_{x_{1:j}}+n_{x_{1:j}}(z_{1:m})$ (19)
and set $\gamma_{x_{1:0}}(z_{1:m},\alpha)=0$.
Theorem 5.1.
Consider the case (10). Then $N\,|\,Z_{1:m}=z_{1:m}$ has a PMF $p(n\,|\,z_{1:m})$ with $p(0\,|\,z_{1:m})\propto p_{0}$ and for every $n\in\mathbb{N}$,
$p(n\,|\,z_{1:m})\propto p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{b}\left(\gamma_{x_{1:j-1},0}(z_{1:m},\alpha),\gamma_{x_{1:j-1},1}(z_{1:m},\alpha)\right)}{{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{n_{x_{1:n}}(z_{1:m})}}}.$ (20)
where the constant of proportionality depends on $z_{1:m}$. Moreover, conditioned on both $N$ and $Z_{1:m}=z_{1:m}$, we have that
$\hskip 35.56593ptY^{(N)}\,|\,N,Z_{1:m}=z_{1:m}\sim\mathrm{FBS}(\gamma^{(N)}(z_{1:m},\alpha^{(N)}))$ (21)
and
$Y^{(>N)}\,|\,N,Z_{1:m}=z_{1:m}\sim\mathrm{BS}(\alpha^{(>N)})$ (22)
are independent.
Remark 11.
The PMF in (20) does not belong to a known parametric family. Since $Y^{(N)}$ is of varying dimension $12^{N}-$ 1, one may suggest to use the reversible jump Markov chain Monte Carlo algorithm [13, 17] for simulating from the posterior density $p(y^{(n)}\,|\,z_{1:m})\propto p(y^{(n)},z_{1:m})$, cf. (B1). However, it is much easier to use Theorem 5.1, since $Y$ conditioned on $(N,Z_{1:m})$ has a conjugate BS-prior, whilst posterior simulation of $N$ is straightforward: First, $N\,|\,Z_{1:m}=z_{1:m}$ can be sampled either exactly, or approximately, or in an asymptotically exact way via a Metropolis-Hastings algorithm as follows. If $N_{\max}$ is an upper bound for $N$, then sample $N$ exactly from a categorical distribution over $\{0,\ldots,N_{\max}\}$ with unnormalized weights given by the right hand side in (20) for $n=0,\ldots,N_{\max}$. This method may also be used to provide an approximate simulation if $N_{\max}$ is a user-defined suitable upper bound for $N$. Alternatively, use a Metropolis-Hastings algorithm with equilibrium density given by (20) (we never found a need for such an algorithm in the examples considered later in this paper). Second, simply use (21) and (22) when simulating $Y$ conditioned on $N$ and $Z_{1:m}=z_{1:m}$.
The posterior mean of the random PDF in (6) is the optimal Bayesian point estimator under the squared-error loss function [34]. The following corollary becomes useful for the calculation of this estimator.
Corollary 5.2.
The posterior mean of the random GFPT1 PDF in (6) is given by
${\mathrm{E}}[f(x\,|\,Y^{(N)})\,|\,Z_{1:m}=z_{1:m}]=\sum_{n=0}^{\infty}\frac{p(n\,|\,z_{1:m})}{\ell_{x_{1:n}}}\prod_{j=1}^{n}\frac{\gamma_{x_{1:j}}(z_{1:m},\alpha)}{\gamma_{x_{1:j-1},0}(z_{1:m},\alpha)+\gamma_{x_{1:j-1},1}(z_{1:m},\alpha)}$ (23)
if $x=.x_{1}x_{2}\ldots\in[0,1)$.
Remark 12.
The right hand side in (23) does not depend on $Y^{(N)}$ and the term $p(n\,|\,z_{1:m})$ may be calculated/approximated as discussed in Remark 11. Hence, the posterior mean in (23) may be determined either by a numerical approximation or by the Monte Carlo estimate of
$\frac{1}{\ell_{x_{1:N}}}\prod_{j=1}^{N}\frac{\gamma_{x_{1:j}}(z_{1:m},\alpha)}{\gamma_{x_{1:j-1},0}(z_{1:m},\alpha)+\gamma_{x_{1:j-1},1}(z_{1:m},\alpha)}$
when $N$ is sampled from $p(n\,|\,z_{1:m})$. We call such an estimate for a density estimate of the posterior GFPT1 PDF (or just density estimate).
5.1.2 Posterior simulation under a GFPT2 prior
Now, consider the posterior distribution of $(Y,R,N)$ for the case (14) where a GFTP2 prior has been specified. The following theorem derives the posterior distribution of first $N\,|\,R$, second $R\,|\,N$, and third $Y\,|\,(N,R)$, where we use the following notation. Recall the definitions (12) and (19). Let $R_{0}=(R_{x_{1:n-1},0}\,|\,x_{1:n-1}\in\{0,1\}^{n-1},n\in\mathbb{N})$. For $j\in\mathbb{N}_{0}$, to stress the dependence of $R^{(j)}_{0}$ (or equivalently of $R^{(j)}$), we use capital letters: For $x_{1:j}\in\{0,1\}^{n}$, let $N_{x_{1:j}}=n_{x_{1:j}}(z_{1:m})$ be the the number of observations $z_{i}$ falling in $I_{x_{1:j}}$, which depends on $R^{(j)}$. Similarly, let $\Gamma_{x_{1:j}}=\gamma_{x_{1:j}}(z_{1:m},\alpha)$ and $L_{x_{1:j}}=\ell_{x_{1:j}}$, which both depend on $R^{(j)}$. For $n\in\mathbb{N}$, $Y^{(n)}$ can be identified by $Y^{(n)}_{0}=(Y_{x_{1:j-1},0}\,|\,x_{1:j-1}\in\{0,1\}^{j-1},j=1,\ldots,n)$ since $Y_{x_{1:j-1},1}=1-Y_{x_{1:j-1},0}$. Let $Y^{(0)}_{0}=Y^{(0)}=\emptyset$. For $n\in\mathbb{N}_{0}$, the dimension of $Y^{(n)}_{0}$ is $12^{n}-$ 1. For $j\in\mathbb{N}$ and $x_{1:j-1}\in\{0,1\}^{j-1}$, let $\nu_{x_{1:j-1}}$ be the Lebesgue measure on $[0,1)$ if $\alpha_{x_{1:j-1},0}>0$ and $\alpha_{x_{1:j-1},1}>0$, and $\nu_{x_{1:j-1}}$ be the Dirac measure concentrated at $k\in\{0,1\}$ if $\alpha_{x_{1:j-1},k}=0$ and $\alpha_{x_{1:j-1},1-k}>0$. For $n\in\mathbb{N}$, let $A_{n}=[0,1)^{2^{n}-1}$ be equipped with the corresponding Borel $\sigma$-algebra $\mathcal{F}_{n}$. Let $\mu_{n}$ be the product measure on $\mathcal{F}_{n}$ given by
$\mu_{n}=\prod_{j=1}^{n}\prod_{{x_{1:j-1}}\in\{0,1\}^{j-1}}\nu_{x_{1:j-1}}.$ (24)
Equip $A_{0}=[0,1)^{0}=\{\emptyset\}$ with the trivial $\sigma$-algebra $\mathcal{F}_{0}$ and let $\mu_{0}=\mathrm{FBS}(\alpha^{(0)})$, cf. Definition 4.3.
Theorem 5.3.
Consider the case (14). Then $N$ conditioned on both $R$ and $Z_{1:m}=z_{1:m}$ has a PMF $p(n\,|\,R,z_{1:m})$ with $p(0\,|\,R,z_{1:m})\propto p_{0}$ and for every $n\in\mathbb{N}$,
$p(n\,|\,R,z_{1:m})\propto p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{b}\big{(}\Gamma_{x_{1:j-1},0},\Gamma_{x_{1:j-1},1}\big{)}}{\prod_{x_{1:n}\in\{0,1\}^{n}}L_{x_{1:n}}^{N_{x_{1:n}}}}$ (25)
where the constant of proportionality depends on $(R,z_{1:m})$. Further, for any $n\in\mathbb{N}_{0}$ with $p_{n}>0$, conditioned on both $N=n$ and $Z_{1:m}=z_{1:m}$, we have that $R^{(n)}$ and $R^{(>n)}$ are independent, where
$\hskip 34.14322ptR^{(>n)}\,|\,N=n,Z_{1:m}=z_{1:m}\sim\mathrm{BS}(\beta^{>(n)})$ (26)
depends only on $n$, and where $R^{(n)}$ or more precisely $R_{0}^{(n)}$ has a density with respect to $\mu_{n}$ which is given by $p(r_{0}^{(n)}\,|\,n,z_{1:m})\propto 1$ when $n=0$ and for every $r_{0}^{(n)}\in(0,1)^{2^{n}-1}$ when $n>0$ by
$p(r_{0}^{(n)}\,|\,n,z_{1:m})\propto\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{b}\big{(}\Gamma_{x_{1:j-1},0},\Gamma_{x_{1:j-1},1}\big{)}r_{x_{1:j-1,0}}^{\beta_{x_{1:j-1,0}}-1}r_{x_{1:j-1,1}}^{\beta_{x_{1:j-1,1}}-1}}{\prod_{x_{1:n}\in\{0,1\}^{n}}L_{x_{1:n}}^{N_{x_{1:n}}}}$ (27)
where the constant of proportionality depends on $(n,z_{1:m})$ and where $L_{x_{1:n}}$ depends on $R_{0}^{(n)}=r_{0}^{(n)}$. Finally, conditioned on both $(R,N)$ and $Z_{1:m}=z_{1:m}$, we have that
$Y^{(N)}\,|\,R,N,Z_{1:m}=z_{1:m}\sim\mathrm{FBS}(\Gamma^{(N)})$ (28)
and
$Y^{(>N)}\,|\,R,N,Z_{1:m}=z_{1:m}\sim\mathrm{BS}(\alpha^{(>N)})$ (29)
are independent, where $\Gamma^{(N)}$ depends only on $R$ through $R_{0}^{(N)}$.
Remark 13.
We use a Metropolis-Hastings within Gibbs sampler which alternates between updating
  1. (I)
    $N\,|\,R,Z_{1:m}=z_{1:m}$,
  2. (II)
    $R\,|\,N,Z_{1:m}=z_{1:m}$,
  3. (III)
    $Y\,|\,R,N,Z_{1:m}=z_{1:m}$.
Steps (I) and (III) are done in the same way as in Remark 11, using (25) for step (I) and (28)–(29) for step (III). In step (II), $R^{(>n)}\,|\,N=n,Z_{1:m}=z_{1:m}$ is simply distributed as in (26) and is independent of $R^{(n)}\,|\,N=n,Z_{1:m}=z_{1:m}$, which follows the unnormalized density (27), where we use a Metropolis-Hastings update.
5.2 Consistency for the posterior of $N$
In this section, we consider the posterior distribution for the case (10) where a GFTP1 prior has been specified but assume $Z_{1},\ldots,Z_{m}$ are i.i.d. according to a PDF $f^{*}$, which we refer to as the true distribution of $Z_{1:m}$. We focus on the asymptotic behaviour of the latent variable $N$ when $m\rightarrow\infty$. At the end of Section 5.2.1 (Remark 15) we motivate the use of GFPT2 priors.
We assume that $f^{*}$ is Lebesgue almost everywhere LSC so that Theorem 3.2 applies. We use the notation $N$ for both the case of the GFPT1 model, with $N$ as in Definition 4.4, and the case of the true model, with $N$ given by the coupling construction in Theorem 3.2 where $f$ is replaced by $f^{*}$ when considering the term $c_{x_{1:n}}$ in (3). This should not cause any confusion, since we denote probabilities calculated with respect to the true distribution by $\mathrm{P}^{*}$, while $\mathrm{P}$ still refers to the distribution under the GFPT1 model. We also use the notation $\mathrm{P}^{*}$ when considering the distribution of the stochastic process $Z_{1},Z_{2},\ldots$ under the true model.
Recall that $p_{n}=\mathrm{P}(N=n)$ is the PMF of $N$ and $p(n\,|\,Z_{1:m})=\mathrm{P}(N=n\,|\,Z_{1:m})$ is the PMF under the posterior distribution of $N$, cf. (20). Let $\mathcal{N}$ be the support of the prior for $N$. For the following Sections 5.2.15.2.2, assume that for some $0\leq n_{\min}\leq n_{\max}\leq\infty$ and some $K>0$,
[…]
The last inequality in (30) is akin of local stability condition, cf. [30].
5.2.1 When $N$ is bounded or not
The following theorems consider consistency of the posterior distribution of $N$ under the true distribution and depending on whether $N$ is bounded (Theorem 5.4) or not (Theorem 5.5).
Theorem 5.4.
Suppose that under the true distribution $n^{*}\in\mathbb{N}_{0}$ digits are sufficient in the sense that
$\mathrm{P}^{*}(N<n^{*})<\mathrm{P}^{*}(N\leq n^{*})=1,$ (31)
cf. Remark 4. In addition, assume $n^{*}\in\mathcal{N}$ and the $\alpha_{x_{1:n}}$ with $n\in\mathbb{N}$ and $x_{1:n}\in\{0,1\}^{n}$ are bounded. Then
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}p(n^{*}\,|\,Z_{1:m})=1\right)=1$ (32)
where $p(n^{*}\,|\,Z_{1:m})$ is given by (20).
Remark 14.
By Theorem 3.2, (31) is equivalent to that $f^{*}$ is constant over all subintervals of the NBP $I$ after level $n^{*}$, that is, for every $x_{1:n^{*}}\in\{0,1\}^{n^{*}}$, $f^{*}$ is constant on $I_{x_{1:n^{*}}}$ and if $n^{*}>0$ then there is some $x_{1:(n^{*}-1)}\in\{0,1\}^{n^{*}-1}$ so that $f^{*}$ is not constant on $I_{x_{1:(n^{*}-1)}}$. We interpret (32) as it is asymptotic consistent to estimate the sufficient number of digits by the posterior distribution of $N$.
Theorem 5.5.
If $\mathrm{P}^{*}(N\leq n)<1$ for all $n\in\mathbb{N}_{0}$, then for any $n\in\mathbb{N}_{0}$,
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}p(n\,|\,Z_{1:m})>0\right){=0.}$ (33)
Remark 15.
The condition in Theorem 5.5 states that under the true distribution, $N$ is unbounded. Hence, (33) establishes asymptotic consistency, since a posteriori $N$ can be as large as it should be as $m$ increases.
The class of PDFs $f^{*}$ for which Theorem 5.5 applies encompasses piecewise constant PDFs over intervals $J_{1},\ldots,J_{K}$ that do not agree with the NBP $I$. That is, $f^{*}(x)=f_{k}$ if $x\in J_{k}$, and there exists at least one $J_{k}$ such that $J_{k}\neq I_{x_{1:n}}$ for every $n\in\mathbb{N}_{0}$ and $x_{1:n}\in\{0,1\}^{n}$.
For example, consider $I$ to be the standard diadic partitions of $[0,1)$ and $f^{*}$ to be the PDF of a $\mathrm{Unif}[0,0.2)$ random variable. Then the condition in Theorem 5.5 is satisfied, so under the GFPT1 prior and as $m\to\infty$, a posteriori $N$ is unbounded. On the other hand, we show numerically in Section 7.2 that under various GFPT2 priors, a posteriori $N$ is bounded (effectively always smaller than 4) even if $m$ is large. Thus, the use of GFPT2 models that adaptively “learn the nested partitions” is appealing.
5.2.2 Round-off errors
We now analyse the impact on posterior inference caused by two types of round-off errors in the data.
Theorem 5.6.
Assume $\mathcal{N}=\mathbb{N}_{0}$ and the data are observed with a precision of $\bar{n}\in\mathbb{N}_{0}$ digits such that for $i=1,\ldots,m$, if $\bar{n}>0$ then for some $(x^{i}_{1},x^{i}_{2},\ldots,x^{i}_{\bar{n}})\in\{0,1\}^{\bar{n}}$,
$z_{i}=.x^{i}_{1}x^{i}_{2}\ldots x^{i}_{\bar{n}}00\ldots,$ (34)
and if $\bar{n}=0$ then $z_{i}=0$. Correspondingly, let $\tilde{Z}_{i}$ be given by the first $\bar{n}$ digits of $Z_{i}$ and by 0’s for the remaining digits. Then, for every $n\in\mathbb{N}_{0}$,
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}p(n\,|\,\tilde{Z}_{1:m})=0\right)=1.$
Theorem 5.6 shows that the round-off error introduced by using 0’s after the precision $\bar{n}$ leads to an inconsistent posterior distribution of $N$ as $m\to\infty$. This is still true if we use another fixed digits rule than in (34), e.g., if we replace the 0’s by 1’s after level $\bar{n}$.
The following corollary follows immediately from Theorems 3.2 and 5.4. For $i=1,\ldots,m$, let $Z_{i}=.X^{i}_{1}X^{i}_{2}\ldots$.
Corollary 5.7.
Assume the round-off error is treated as being uniformly distributed after level $\bar{n}\in\mathbb{N}_{0}$, that is, under the GFPT1 prior we have $N\leq\bar{n}$ and
.$X^{i}_{\bar{n}+1}X^{i}_{\bar{n}+2}\ldots\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\mathrm{Unif}[0,1)\quad\mbox{for }i=1,\ldots,m.$ (35)
Further, assume that under the true distribution, in the sense of (31), $n^{*}\leq\bar{n}$ digits are sufficient where $n^{*}\in\mathcal{N}$, and that all the $\alpha_{x_{1:n}}$ are bounded. Then
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}p(n^{*}\,|\,Z_{1:m})=1\right)=1.$ (36)
Remark 16.
Under the GFPT1 prior, $N\leq\bar{n}$ implies (35), cf. (8). Equation (36) shows it is asymptotic consistent to estimate the sufficient number of digits by the posterior distribution of $N$.
Under the true distribution, if $n^{*}$ digits are sufficient, then $n^{*}$ is the largest possible value of $N$. Ideally we should therefore have that $n^{*}\in\mathcal{N}\subseteq\{0,\ldots,\bar{n}\}$ but in practice we do not know the value of $n^{*}$ (provided it exists). This suggests to let $\mathcal{N}=\{0,\ldots,\bar{n}\}$. This is also an intuitive suggestion as the data contains no information beyond level $n^{*}$ of the nested binary partitions.
5.3 Consistency for the posterior of $f$
In this section we assume the same setting as in Section 5.2 and consider the random GFPT1 PDF $f(\cdot\,|\,Y^{(N)})$ and the random GFPT2 PDF $f(\cdot\,|\,Y^{(N)},R^{(N)})$ given by (6) and (13), respectively. For $i=1,2$, using the short hand notation $f=f(\cdot\,|\,Y^{(N)})$ if $i=1$ and $f=f(\cdot\,|\,Y^{(N)},R^{(N)})$ if $i=2$, recall that a posteriori $f\,|\,Z_{1:m}\sim\Pi_{i}(\cdot\,|\,Z_{1:m})$, cf. (11) and (15). As the sample size increases, we will show that $\Pi_{i}(\cdot\,|\,Z_{1:m})$ concentrates on infinitesimal small neighbourhoods of $f^{*}$ with respect to the Kolmogorov-Smirnov distance $d_{\mathrm{KS}}$ and the Hellinger distance $d_{\mathrm{H}}$: For two PDFs $f_{1}$ and $f_{2}$ on $[0,1)$, if $F_{i}(x)=\int_{0}^{x}f_{i}(x)\,\mathrm{d}x$ for $i=1,2$ and $10\leq x<$ 1, recall that
$d_{\mathrm{KS}}(f_{1},f_{2})=\sup_{x\in[0,1)}|F_{1}(x)-F_{2}(x)|$
and
$d_{\mathrm{H}}(f_{1},f_{2})=\frac{1}{2}\left(\int_{0}^{1}(\sqrt{f_{1}(x)}-\sqrt{f_{2}(x)})^{2}\mathrm{d}x\right)^{1/2}.$
Theorem 5.8.
Suppose that $\int_{0}^{1}f^{*}(x)\log f^{*}(x)\,\mathrm{d}x<\infty$, $p_{n}>0$ for all $n\in\mathbb{N}_{0}$, $\alpha_{x_{1:n}}>0$ and, in case of GFPT2 priors, $\beta_{x_{1:n}}>0$ whenever $n\in\mathbb{N}$ and $x_{1:n}\in\{0,1\}^{n}$. Then, for any $\varepsilon>0$ and $i=1,2$,
$\mathrm{P^{*}}\left(\lim_{m\rightarrow\infty}\Pi_{i}(d_{\mathrm{KS}}(f,f^{*})<\varepsilon\,|\,Z_{1:m})=1\right)=1$
where $f=f(\cdot\,|\,Y^{(N)})$ if $i=1$ and $f=f(\cdot\,|\,Y^{(N)},R^{(N)})$ if $i=2$.
Theorem 5.8 gives sufficient conditions for the posterior of $f$ to concentrate on vanishing neighbourhoods of $f^{*}$ in the weak topology, since convergence with respect to $d_{\mathrm{KS}}$ is stronger than weak covergence of probability measures on Euclidean spaces (see e.g. Theorem 6 in [15]). However, as discussed in [14], consistency under the weak topology is often considered “too weak” for density estimation purposes. Instead the next theorem gives sufficient conditions for consistency with respect to the Hellinger distance. Convergence with respect to $d_{\mathrm{H}}$ is stronger than weak convergence, and the topology induced by the Hellinger distance is equivalent to the one induced by the total variation norm (see e.g. [24]).
Theorem 5.9.
In addition to the assumptions of Theorem 5.8, suppose $p_{n}<Ce^{-cn}$ for some constants $C>0$ and $c>0$ and for all $n\in\mathbb{N}$. Then, for any $\varepsilon>0$,
$\mathrm{P^{*}}\left(\lim_{m\rightarrow\infty}\Pi_{1}\left(d_{\mathrm{H}}(f(\cdot\,|\,Y^{(N)}),f^{*})<\varepsilon\,|\,Z_{1:m}\right)=1\right)=1.$
Remark 17.
For instance, the condition $0<p_{n}<Ce^{-ck}$ is satisfied by the Poisson distribution. By contrast, posterior consistency for standard Pólya tree priors was established in [2] (see also Theorem 7.16 in [14]) under the assumptions that the parameters $\alpha_{x_{1:n}}$ grow exponentially fast with $n$. As noted in [14] this assumption is too strong, since practitioners usually follow [25] by setting $\alpha_{x_{1:n}}=\alpha_{0}n^{2}$. See also [7, 16] for further developments.
6 Modelling multi-scale human data via Benford’s law and generalized Pólya trees
Real-world datasets – such as those involving incomes, city sizes, or seismic magnitudes – often span multiple orders of magnitude [8]. Modelling such data poses a challenge for Bayesian density estimation because traditional methods may have difficulty capturing both the coarse-scale variability and the fine-scale structure [35]. We show here how a simple extension of a GFPT1 prior can be used to obtain a simple, yet powerful, Bayesian model for density estimation in scenarios where data span multiple orders of magnitudes and their digits exhibit the Newcomb–Benford law. The main idea lies in independently modelling the order of magnitude of the data and their digits, by assuming a GFPT1 prior for the digits.
[…]
Consider data $Z_{i}>0$, $i=1,\ldots,m$, represented in terms of their order of magnitudes and their base-$q$ digits as $Z_{i}=(M_{i},\tilde{Z}_{i})$, such that $Z_{i}=q^{M_{i}+1}\tilde{Z}_{i}$ (see again Section 4.3). Assume
$\tilde{Z}_{1},\ldots,\tilde{Z}_{m}\,|\,Y,N\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}f\big{(}\cdot\,|\,Y^{(N)}\big{)}\sim\mathrm{BPT}_{q}.$
Further, assume the $M_{i}$’s take values in $T\subset\mathbb{Z}$ of cardinality $|T|<\infty$ such that
$\mathrm{P}(M_{i}=k\,|\,\omega)=\omega_{k},\quad k\in T,$
where we impose the prior
$\omega=(\omega_{k}\,|\,k\in T)\sim\mathrm{Dir}(\eta)$
where $\mathrm{Dir}(\eta)$ denotes the Dirichlet distribution with parameter $\eta\in\mathbb{R}_{+}^{|T|}$. Finally, assume conditioned on $(Y,N,\omega)$ that the $M_{i}$’s and the $\tilde{Z}_{i}$’s are independent, and a priori that $\omega$ is independent of $(Y,N)$. We call the distribution of $(Y,N,\omega)$ a base-$q$ multiscale Benford Pólya tree ($\mathrm{MBPT}_{q}$) prior.
The $\mathrm{MBPT}_{q}$ prior induces a scale invariant (in the terminology of [22]) random density on $T\times[0,1)$ equipped with the product of counting measure on $T$ and Lebesgue measure on $[0,1)$. This random density is given by
$f(x\,|\,Y^{(N)},\omega)=\omega_{k}\,q^{N}\prod_{j=1}^{N}\prod_{d=0}^{q}Y^{I(\tilde{x}_{j}=d)}_{\tilde{x}_{1:j-1},d}\quad\text{if }x=(k,\tilde{x})\in T\times[0,1)$
where $I[\cdot]$ denotes the indicator function. We call it the random $\mathrm{MBPT}_{q}$ PDF. When considering its posterior distribution, we call it the posterior $\mathrm{MBPT}_{q}$ PDF.
Posterior computation is straightforward. Indeed, $(Y,N)$ and $\omega$ are independent also a posteriori, with the posterior of $(Y,N)\,|\,Z_{1:m}$ given by a slight modification of Theorem 5.1. Moreover, $\omega\,|\,Z_{1:m}\sim\mathrm{Dir}(\eta^{p})$ where $\eta^{p}_{k}=\eta_{k}+\sum_{i=1}^{m}I[M_{i}=k]$ for $k\in T$.
7 Numerical illustrations
This section investigates several aspects of our models via numerical simulations. Prior elicitation follows the discussion in Section 4.3. Unless otherwise specified, a priori $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,20\}$ where the Poisson distribution has mean 5.
We need the following notation. For two probability densities $f$ and $g$, $\mathrm{TV}(f,g)=\frac{1}{2}\int|f(x)-g(x)|\,\mathrm{d}x$ denotes their total variation distance. For a curve $g:[a,b]\rightarrow\mathbb{R}$, denote its length by $L_{g}$ and let $W_{g}=L_{g}-|b-a|$. We use $W_{g}$ to compare the “wigglyness” of two PDFs: Let $f_{0}$ be a reference PDF, $\hat{f}_{j}$, $j=1,2$ be two estimators of $f_{0}$, and $W_{j}=W_{\hat{f}_{j}-f_{0}}$ for $j=1,2$. Then we say that $\hat{f}_{1}$ is less wiggly than $\hat{f}_{2}$ if $W_{1}<W_{2}$. Clearly, for any $\hat{f}$, $W_{\hat{f}-f_{0}}\geq 0$ and we have $W_{\hat{f}-f_{0}}=0$ if and only if $\hat{f}=f_{0}$ (Lebesgue almost everywhere).
7.1 Posterior results when using GFPT1 and Pólya tree priors
We consider three data generating processes:
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\mathrm{Unif}(0,0.5)$ (DG1)
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{6}\mathrm{Unif}(0,0.25)+\frac{1}{2}\mathrm{Unif}(0.125,0.25)+\frac{1}{3}\mathrm{Unif}(0.5,1)$ (DG2)
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\mathcal{N}(0.5,0.1)|_{[0,1)}$ (DG3)
[…]
Figure 1 shows the posterior density estimates (see Remark 12) for the 12 cases. For all datasets, the PT prior yields extremely wiggly posterior density estimates when $\alpha_{0}=0.05$ (dotted orange line), and these get only partially less wiggly for $\alpha_{0}=2$ (dotted blue line). On the other hand, for datasets DG1 and DG2 (the two first columns), the GFPT1 prior yields precise posterior density estimates for both values of $\alpha_{0}$ (the solid blue and orange lines). However, for DG3 and $\alpha_{0}=2.0$, the posterior density estimate under the GFPT1 prior is not very precise (last column, solid blue line). This happens because the posterior of $N$ concentrates on small values leading to a coarse approximation. When $\alpha_{0}=0.05$ instead, the posterior of $N$ concentrates to higher values leading to a more accurate density estimate (last column, solid orange line).
Figure 1: Posterior results for first simulation study in Section 7.1: Density estimates under GFPT1 (solid lines) and PT priors (dotted lines) when $\alpha_{0}=2$ (blue lines) and $\alpha_{0}=0.05$ (orange lines) where the three columns correspond to the data examples DG1–DG3, respectively.
In our second simulation study, we confirmed such insights by generating $100$ independent datasets from each data generating process with
$m\in\{50,100,1000,5000,10000\}$
and where each dataset was fitted using PT and GFPT1 priors with
$\alpha_{0}\in\{0.05,0.1,2.0,10.0\}.$
Figures 24 show the mean across the 100 independent datasets of
  1. (i)
    the total variation distance distance between the density of the true data generating process and the posterior density estimate (for short the true and estimated densities),
  2. (ii)
    the measure of wigglyness of the difference between the true and estimated densities,
  3. (iii)
    the a posteriori expectation of $N$,
respectively. As expected Figures 24 show that larger values of $\alpha_{0}$ correspond to smaller values of $N$ (a posteriori) and vice versa. This is particularly evident for datasets DG1 and DG3. In general, we see that GFPT1 priors outperform PT priors in terms of the quality of the posterior density estimate. Moreover, larger $\alpha_{0}$ clearly yield less wiggly functions when using PT priors, while this dependence is not so clear when using GFPT1 priors. In conclusion, we find that $\alpha_{0}=0.1$ provides the best trade off between wigglyness and accuracy of the density estimate, and we suggest using it as default value for GFPT1 priors.
Figure 2: Posterior results for the second simulation study in Section 7.1: Total variation (TV) distance between true and estimated densities as a function of the sample size $m$ where the three columns correspond to the data examples DG1–DG3, respectively. Figures are averaged over 100 independent replicates. Solid and dashed lines refer to the use of GFPT1 and PT priors, respectively. Different colors refer to different values of $\alpha_{0}$.
Figure 3: Posterior results for the second simulation study in Section 7.1: $W$-function for the difference between true and estimated densities. All details are as in the caption of Figure 2.
Figure 4: Posterior results for the second simulation study in Section 7.1: Posterior expectation of $N$ for the simulation in Section 7.1. All details are as in the caption of Figure 2.
7.2 Posterior results when using GFPT1 and GFPT2 priors
This section investigates through simulation studies if there is any gain in extending GFPT1 models to GFPT2 models. For this we consider six different data generating processes, namely DG1–DG3 as in Section 7.1 and
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\mathrm{Unif}[0,0.2)$ (DG4)
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{2}\mathrm{Unif}[0,0.2)+\frac{1}{2}\mathrm{Unif}[0.7,0.9)$ (DG5)
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{2}\mathrm{B}(2,15)+\frac{1}{2}\mathrm{B}(15,2).$ (DG6)
We consider a GFPT1 prior where $\alpha_{0}=0.1$ (as suggested at the end of Section 7.1) and a GFPT2 prior where $\alpha_{x_{1:n}}=\alpha_{0}n^{2}$ with $\alpha_{0}\in\{0.1,2.0\}$ (in Section 7.1, these values of $\alpha_{0}$ provided the best GFPT1 posterior estimates). As far as the prior for $R$ is concerned (cf. Defintion 4.5), we assume $\beta_{x_{1:n}}=\beta_{0}$ for all $n\in\mathbb{N}$ and all $x_{1:n}\in\{0,1\}^{n}$, so that the prior mean for the NBP $I$ is given by the standard dyadic partitions employed by the GFPT1 model. We let $\beta_{0}\in\{0.5,2.0,5.0\}$, which allowed us to understand the sensitivity of GFPT2 priors: When $\beta_{0}=0$, the $\mathrm{BS}(\beta)$ prior for $R$ assigns significant mass to values of $R_{x_{1:n}}$ near to 00 or $1$, corresponding to almost empty intervals. Instead when $\beta_{0}\in\{2.0,5.0\}$, the $\mathrm{BS}(\beta)$ prior for $R$ causes a random NBP which is more and more concentrated around the standard dyadic partitions. We also tried increasing the value of of $\beta_{0}$ when increasing the level $n$ of the partitions, but this led to worse mixing of the MCMC algorithm for posterior simulations (see Remark 13; we used 10,000 iterations and discarding the first 1,000 as burn-in). In our examples, running this MCMC algorithm on a standard laptop takes only around five seconds.
Figure 5: Total variation distance (TV) between true and estimated densities (first two rows of plots) and posterior expectation of $N$ (last two rows). Each boxplot summarizes the results from 100 independent replicates of the indicated data generating model. The $\beta_{0}$ parameter of the GPFT2 model varies across the $x$-axis of each plot.
For each data generating process, we simulated 1,000 observations and repeated each analysis on 100 independently simulated datasets. Figure 5 shows for the different data generating processes and parameter values, the total variation distance between true and estimated densities as well as the posterior mean of $N$. We observer the following. First, the parameter $\beta_{0}$ does not seem to play a significant role in posterior inference. Second, when using GFPT2 priors, as in the case of GFPT1 priors, higher values of $\alpha_{0}$ correspond to smaller values of $N$ a posteriori. In some scenarios (e.g., DG1 and DG4) this yields also a more accurate estimation of the density because the density needs only a coarse partition to be well approximated. In other scenarios (e.g., DG3 and DG6), this results in a much poorer density estimate. In both such cases, the data generating density entails that the number of sufficient digits is infinite, and a finer partition results in better density estimates. Focusing specifically on DG4 and DG5, observe that those data generating densities are piecewise continuous on intervals that do not agree with the standard diadic partitions employed for the GFPT1 prior. As a result, under the GFPT1 prior, a posterior $N$ tends to be large in those scenarios, while especially when $\alpha_{0}=2$ the GFPT2 model correctly captures the partitions needed to approximate those densities with a small number of intervals, resulting in smaller $N$ a posteriori. In particular, for DG4, using the GFPT2 prior with $\alpha_{0}=2$ achieves almost a perfect density estimate.
7.3 Fitted models for real datasets spanning multiple orders of magnitude
We consider here five real-world datasets:
  1. (i)
    Twitter records the number of friends of 40,000 users,
  2. (ii)
    eurodist contains the pairwise distances (in kilometers) between 21 cities in Europe,
  3. (iii)
    GDP collects the gross domestic products (in dollars) of 196 countries,
  4. (iv)
    census is the number of citizens in more than 19,000 US cities as of 2009,
  5. (v)
    income records the personal income of more than 50,000 inhabitants in California as of 2023.
Data for the first four dataset can be found at https://github.com/jasonlong/benfords-law, while data for the income dataset is publicly available from https://www.census.gov/data.html. Table 1 reports the empirical range (maximum and minimum value of the data) in the different datasets. The datasets in Table 1 span four, three, eight, six, and seven orders of magnitude, respectively, making them particularly well suited to test the adequacy of the base-2 and base-10 multiscale Benford Pólya tree priors.
Dataset Twitter Eurodist GDP Census Income
Range $(101,2.3\cdot 10^{5})$ $(158,4532)$ $(5\cdot 10^{7},10^{14})$ $(1,8\cdot 10^{6})$ $(4,1.3\cdot 10^{7})$
Table 1: Empirical range of the observations in the different datasets analyzed in Section 7.3.
We compare posterior results when using five different priors, namely $\mathrm{MBPT}_{2}$ and $\mathrm{MBPT}_{10}$ priors, GFPT1 and GFPT2 priors fitted on max-scaled data (i.e., for each dataset, we divide all observations by the maximum value), and a Dirichlet process mixture (DPM) of Gaussian densities [11, 14, 28, 32], which is the de-facto standard model for Bayesian density estimation. The parameters of the GFPT1 and GFPT2 priors are selected following Section 4.3 and the insights developed in Section 7.1. Specifically, $\alpha_{x_{1:n}}=0.1n^{2}$ for both priors, and $\beta_{x_{1:n}}=02$ for the CFPT2 prior. For the prior on $N$, we proceed as follows. Let $n_{\max}$ be the maximum number of base-10 digits recorded in a given dataset. Then, for the $\mathrm{MBPT}_{10}$ prior, assume that $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,n_{\max}\}$, where the Poisson distribution has mean $n_{\max}/2$. For the $\mathrm{MBPT}_{2}$, GFPT1, and the GFPT2 priors, assume that $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,\lceil\log_{2}(10^{n_{\max}})\rceil\}$, where the Poisson distribution has mean $\log_{2}(10^{n_{\max}})/2$. This ensures that under all five priors of $N$, the number of possible values of the sufficient digits is approximately the same. Further, for the $\mathrm{MBPT}_{2}$ and $\mathrm{MBPT}_{10}$ priors, we fix $T$ as the span of the orders of magnitudes of the data, set $\omega_{k}\propto 1$ for all $k$, and let $c_{n}$ from (18) be given by $c_{n}=c_{0}n^{2}$ for $n\in\mathbb{N}$, where $c_{0}=0.1$ in case of $\mathrm{MBPT}_{10}$ and $c_{0}=2.0$ in case of $\mathrm{MBPT}_{2}$. This choice of $c_{0}$ entails that the variance of the random PDF in (6) is approximately equal under both MBPT priors. Finally, for the DPM prior, we use the implementation and follow the default prior elicitation strategy in the BayesMix library [3].
The posterior inferences using the different priors are compared in term of the widely applicable information criterion (WAIC), see [37], which is a consistent estimator of the out-of-sample error [36]. Since the absolute value of the WAIC is irrelevant, we fix the DPM as a baseline prior and report in Table 2 the relative improvement of the WAIC for any model over the DPM (larger values of the relative improvement correspond to better model performance). For datasets eurodist, GDP, and census, all priors seem to perform similarly. For the datasets Twitter and income datasets, the multiscale Benford Pólya tree priors clearly outperforms all other priors, where in particular, the $\mathrm{MBPT}_{10}$ prior shows a remarkable improvement of rougly 19% and 25% over the DPM prior, respectively. Moreover, in all scenarios, the WAIC when using the GFPT2 prior slightly improves as compared to the GFPT1 prior.
Dataset GFPT1 GFPT2 $\mathrm{MBPT}_{2}$ $\mathrm{MBPT}_{10}$ DPM
Twitter -0.015 0.006 0.150 0.189 0.000
Eurodist -0.000 0.007 -0.004 -0.059 0.000
GDP -0.002 0.060 0.059 0.049 0.000
Census -0.031 0.035 0.036 0.084 0.000
Income 0.002 0.045 0.210 0.247 0.000
Table 2: Relative improvement of the WAIC over the Dirichlet process mixture prior.
[…]
$\mathrm{E}[N\,|\,Z_{1:M}]$ $\mathrm{E}[q^{N}\,|\,Z_{1:M}]$
Dataset GFPT1 GFPT2 $\mathrm{MBPT}_{2}$ $\mathrm{MBPT}_{10}$ GFPT1 GFPT2 $\mathrm{MBPT}_{2}$ $\mathrm{MBPT}_{10}$
Twitter 9.0 7.0 12.0 4.0 512.80 128 4096 10000
Eurodist 2.93 2.78 2.25 1.9 7.72 7.51 6.25 98.72
GDP 5.0 5.0 2.92 1.9 32.32 32.00 8.86 99.93
Census 10.0 6.0 4 1.9 1024 64 16 99.99
Income 10.0 6.0 12.0 4.0 1024 64 4096 10000
Table 3: For the five datasets and four priors, posterior expectation of $N$ and $q^{N}$ with $q=10$ for the $\mathrm{MBPT}_{10}$ prior and $q=2$ for the remaining ones.
8 Concluding remarks and open problems
This paper develops a family of generalised finite Pólya tree (GFPT) priors for a random PDF in which the random sufficient digits are modelled directly. Working with digits rather than with more classical quantities such as the mean or the variance is admittedly non-standard and, at first sight, conceptually harder. However, the vast literature developed for prior elicitation in standard Pólya trees can be clearly adapted in our setting as well. As shown by our construction of the multiscale Benford Pólya tree prior, there are situations in which reasoning about the distribution of the digits is both more natural and leads to significant performance gain in the quality of posterior inference.
A natural extensions is to consider multivariate data, either by defining a product nested binary partition on the unit cube, or by relying on coordinate values given by space-filling curves such as Morton’s or Hilbert’s curves. Another idea which we leave for future research is the slight modification of the GFPT2 model obtained by the mixture distribution
$R_{x_{1:n}}\,|\,R_{x_{1:n-1}}\sim\begin{cases}\pi\delta_{\{1\}}+(1-\pi)B(\beta_{x_{1:n-1},0},\beta_{x_{1:n-1},1})&\text{if }R_{x_{1:n-1}}<1\\ \delta_{\{1\}}&\text{if }R_{x_{1:n-1}}=1\end{cases}$
where $10<\pi<$ 1 is a parameter. In this way, we introduce an “optional” stopping of the partitioning in some regions of the domain, similarly to what is done in [38].
Theorem 5.9 establishes under weak assumptions almost sure consistency in the Hellinger distance of the random PDF a posteriori when using GFPT1 priors. We leave it as an open problem to establish similar convergence results under GFPT2 priors.
Ongoing work by us extends the setting in this paper, utilizing the numeric system induced by a continued fraction representation. That is, $X_{1},X_{2},...\in\mathbb{N}$ are the random digits of $X$ such that $X=1/(X_{1}+(1/X_{2}+...))$. Thus, the digits take value in an unbounded space, which complicates inference, yet promises sharper inference because continued fractions give the best rational approximations of real numbers.
Acknowledgements
Supported by The Danish Council for Independent Research — Natural Sciences, grant DFF – 10.46540/2032-00005B.
Supplemental material for:
“Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees”
Appendix A Proofs for Section 3
Proof of Theorem 3.2.
The existence and coupling of $N$ with $X$ such that (3) is satisfied follows from [29].
Let the situation be as in the second part of the theorem. Clearly, $g$ is a PDF on $H$. Recall that a set $U\subseteq H$ is open (with respect to $H$) if it is the intersection of $H$ with an open subset of $\mathbb{R}$. The following facts are well-known, see e.g. [10].
  1. (a)
    A real function $h$ is LSC on $H$ if and only if for every $u>0$ the set $U=\{x\in H\,|\,h(x)>u\}$ is open (with respect to $H$), i.e., for any $x\in U$ there is a neighbourhood $U_{x}$ which means that $x\in U_{x}\subseteq U$ and $U_{x}$ is open (with respect to $H$).
  2. (b)
    A finite sum of LSC functions on $H$ is a LSC function on $H$.
Now, suppose $x\in H$ and $0<u<g(x)$. Since $g(x)=\sum_{n=0}^{\infty}g_{n}(x)$, there exists some $k_{x}\in\mathbb{N}_{0}$ so that $\sum_{n=0}^{k_{x}}g_{n}(x)>u$. By (b), $\sum_{n=0}^{k_{x}}g_{n}$ is a LSC function on $H$, and hence by (a), the set $U_{x}=\{y\in H\,|\,\sum_{n=0}^{k_{x}}g_{n}(y)>u\}$ is open (with respect to $H$). Since $x\in U_{x}\subseteq U$, it follows that $g$ is LSC on $H$. The remaining statements in Theorem 3.2 are obviously true. ∎
Appendix B Proofs for Section 5.1
Proof of Theorem 5.1.
Recall the definition in (24) of the measure $\mu_{n}$ and the corresponding notation $Y_{0}^{(n)},A_{n},\mathcal{F}_{n}$. Then, $Y^{(N)}_{0}$ has state space $A=\cup_{n=0}^{\infty}A_{n}$ which we equip with the smallest $\sigma$-algebra $\mathcal{F}$ which contains $\cup_{n=0}^{\infty}\mathcal{F}_{n}$. Furthermore, define a reference measure $\mu$ on $(A,\mathcal{F})$ so that $\mu(F)=\mu_{n}(F)$ whenever $F\in\mathcal{F}_{n}$ and $n\in\mathbb{N}_{0}$.
Let $n\in\mathbb{N}_{0}$, $y^{(n)}=0$ if $n=0$, and $y^{(n)}=(y_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},j=1,\ldots,n)$ if $n>0$ where $0\leq y_{x_{1:j-1},0}\leq 1$ and $y_{x_{1:j-1},1}=1-y_{x_{1:j-1},0}$. Then we can identity $y^{(n)}$ by $y^{(n)}_{0}=(y_{x_{1:j-1},0}\,|\,x_{1:j-1}\in\{0,1\}^{j-1},1\leq j\leq n)$. Definitions 4.1 and 4.3 give that the distribution of $Y^{(N)}$ (or more precisely $Y^{(N)}_{0}$) is absolutely continuous with respect to $\mu$, with density
$p(y^{(n)})=\begin{cases}p_{0}&\text{if }n=0,\\ p_{n}\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{B}({y_{x_{1:j-1},0}}\,|\,\alpha_{x_{1:j-1,0}},\alpha_{x_{1:j-1,1}})&\text{if }n>0.\end{cases}$
By Remark 8, the distribution of $(Y^{(N)},Z_{1:m})$ is absolutely continuous with respect to the product measure of $\mu$ and the Lebesgue measure on $(0,1)^{m}$, with density $p(z_{1:m},y^{(n)})$ such that for $n>0$,
$p(y^{(n)},z_{1:m})=p(y^{(n)})\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}y_{x_{1:j-1},0}^{n_{x_{1:j-1},0}(z_{1:m})}y_{x_{1:j-1},1}^{n_{x_{1:j-1},1}(z_{1:m})}}{{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{n_{x_{1:n}}(z_{1:m})}}}$ (B1)
$=p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}y_{x_{1:j-1},0}^{n_{x_{1:j-1},0}(z_{1:m})}y_{x_{1:j-1},1}^{n_{x_{1:j-1},1}(z_{1:m})}\mathrm{B}({y_{x_{1:j-1},0}}\,|\,\alpha_{x_{1:j-1},0},\alpha_{x_{1:j-1},1})}{{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{n_{x_{1:n}}(z_{1:m})}}}$
whilst if $n=0$ then $p(y^{(n)},z_{1:m})=p_{0}$.
The conditional distribution of $N$ given $Z_{1:m}=z_{1:m}$ has a PMF $p(n\,|\,z_{1:m})$ with $p(0\,|\,z_{1:m})\propto p(0)$ and where for every $n\in\mathbb{N}$, (B1) gives
$p(n\,|\,z_{1:m})\propto p(n,z_{1:m})=\int_{A_{n}}p(y^{(n)},z_{1:m})\,\mathrm{d}\mu_{n}(y^{(n)}_{0})$
which reduces to (20). It also follows from (B1) that the conditional distribution of $Y^{(N)}$ given $(N,Z_{1:m})=(n,z_{1:m})$ is absolutely continuous with respect to $\mu_{n}$, with a density $p(y\,|\,n,z_{1:m})$ for $y\in A_{n}$ so that if $n>0$,
$p(y^{(n)}\,|\,n,z_{1:m})$
$\propto\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}y_{x_{1:j-1,0}}^{n_{x_{1:j-1,0}}(z_{1:m})}y_{x_{1:j-1,1}}^{n_{x_{1:j-1,1}}(z_{1:m})}\mathrm{B}({y_{1:j-1,0}}\,|\,\alpha_{x_{1:j-1,0}},\alpha_{x_{1:j-1,1}})$
whilst if $n=0$ then $Y^{(N)}=0$. This together with Definition 4.4 and the fact that $X$ and $Y^{(>N)}$ conditioned on $N$ are independent gives the last statement of the theorem. ∎
Proof of Corollary 5.2.
We have
${\mathrm{E}}[f(x\,|\,Y^{(N)},I^{(N)})\,|\,Z_{1:m}=z_{1:m}]={\mathrm{E}}[{\mathrm{E}}[f(x\,|\,Y^{(N)},I^{(N)})\,|\,N,Z_{1:m}]\,|\,Z_{1:m}=z_{1:m}]$
$=\sum_{n=0}^{\infty}p(n\,|\,z_{1:m}){\mathrm{E}}[f(x\,|\,Y^{(N)},I^{(N)})\,|\,N=n,Z_{1:m}=z_{1:m}]$
$=\sum_{n=0}^{\infty}p(n\,|\,z_{1:m}){\mathrm{E}}\left[\frac{\prod_{j=1}^{n}Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}}{\ell_{x_{1:n}}}\,\bigg{|}\,N=n,Z_{1:m}=z_{1:m}\right]$
where the first equality follows from the law of total expectation and because of the conditional independence in (21) and (22), the second equality is based on the law of total probability, and the last equality follows from (6). Thereby, using (21) we obtain (23). ∎
Proof of Theorem 5.3.
In a similar way as the proof of the first statement of Theorem 5.1, it is possible to analytically marginalize out $Y$ from the posterior of $(N,Y,R)$, whereby the two first statements of Theorem 5.3 follow. The last statement follows immediately from Theorem 5.1. ∎
Appendix C Proofs for Section 5.2
Proof of Theorem 5.4.
Theorem 5.4 is obviously true if $\mathcal{N}=\{n^{*}\}$, so assume $\mathcal{N}$ has cardinality at least two.
To get rid of the associated normalizing constant in (20), we will prove the following equivalent statement of (32): For any $\tilde{n}\in\mathcal{N}\setminus\{n^{*}\}$,
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}{p(n^{*}\,|\,Z_{1:m})}/{p(\tilde{n}\,|\,Z_{1:m})}=\infty\right)=1.$ (C2)
To establish (C2) we use the following facts, considering any $n\in\mathcal{N}$ such that $n+1\in\mathcal{N}$. By the law of large numbers,
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}N_{x_{1:n}}/m=\textstyle\int_{I_{x_{1:n}}}f^{*}(v)\,\mathrm{d}v\right)=1$ (C3)
where for any $j\in\mathbb{N}_{0}$ and $x_{1:j}\in\{0,1\}^{j}$, we use $N_{x_{1:j}}=n_{x_{1:j}}(Z_{1:m})$ as a shorthand notation (this should not be confused with the notation used in Section 5.1.2). By (20),
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}=\frac{p_{n}}{p_{n+1}}\times\frac{\prod_{x_{1:(n+1)}\in\{0,1\}^{n+1}}\ell_{x_{1:(n+1)}}^{N_{x_{1:(n+1)}}}}{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{N_{x_{1:n}}}\mathrm{b}(\alpha_{x_{1:n},0}+N_{x_{1:n},0},\alpha_{x_{1:n},1}+N_{x_{1:n},1})}.$
Since $N_{x_{1:n}}=N_{x_{1:n},0}+N_{x_{1:n},1}$ and $\ell_{x_{1:n}}=\ell_{x_{1:n},0}+\ell_{x_{1:n},1}$, we get
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}$ $=\frac{p_{n}}{p_{n+1}}\prod_{x_{1:n}\in\{0,1\}^{n}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{(\ell_{x_{1:n},0}+\ell_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}}}$
×$\hskip 56.9055pt\times\frac{1}{\mathrm{b}(\alpha_{x_{1:n},0}+N_{x_{1:n},0},\alpha_{x_{1:n},1}+N_{x_{1:n},1})}.$ (C4)
Define
$i_{x_{1:n}}=\begin{cases}1&\text{if }\int_{I_{x_{1:n}}}f^{*}(v)\,\mathrm{d}v>0,\\ 0&\text{otherwise}.\end{cases}$
If $i_{x_{1:n}}=0$ then $\mathrm{P}^{*}(N_{x_{1:n}}=0)=1$. For two sequences $a_{1},a_{2},\ldots$ and $b_{1},b_{2},\ldots$ of real numbers, write $a_{m}\asymp b_{m}$ if $\lim_{m\rightarrow\infty}a_{m}/b_{m}=1$. In the remainder of this proof it is implicit that any convergence result hold almost surely under the true distribution as $m\to\infty$. Thus,
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}\asymp{\frac{p_{n}}{p_{n+1}}}\times$
$\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{(\ell_{x_{1:n},0}+\ell_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}$ (C5)
thanks to (C3), (C4), the condition on the $\alpha_{x_{1:n}}$’s, and Stirling’s approximation of the beta function: $\mathrm{b}(x,y)=\sqrt{2\pi}\frac{x^{x-1/2}y^{y-1/2}}{(x+y)^{x+y-1/2}}\left(1+O(1/x)+O(1/y)\right)$.
Now, consider any integer $n\in\mathcal{N}$ with $n>n^{*}$. By Theorem 3.2, for any $y\in I_{x_{1:n}}$ we have that $f^{*}(y)$ is constant and $\int_{I_{x_{1:n}}}f^{*}(v)\,\mathrm{d}v=f^{*}(y)\ell_{x_{1:n}}$. Hence, for $k=0,1$ and $i_{x_{1:n}}=1$, it follows from (C3) that
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}N_{x_{1:n},k}/N_{x_{1:n}}=\ell_{x_{1:n},k}/\ell_{x_{1:n}}\right)=1.$ (C6)
By (C5) and (C6),
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}$ $\asymp\frac{p_{n}}{p_{n+1}}\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}}{(\ell_{x_{1:n},0}+\ell_{x_{1:n},1})^{N_{x_{1:n},0}}}\frac{\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{(\ell_{x_{1:n},0}+\ell_{x_{1:n},1})^{N_{x_{1:n},1}}}$
×$\times\left(\frac{N_{x_{1:n},0}+N_{x_{1:n},1}}{N_{x_{1:n},0}}\right)^{N_{x_{1:n},0}-1/2}\left(\frac{N_{x_{1:n},0}+N_{x_{1:n},1}}{N_{x_{1:n},1}}\right)^{N_{x_{1:n},1}-1/2}$
×$\times[(N_{x_{1:n},0}+N_{x_{1:n},1})^{1/2}/{\sqrt{2\pi}}]$
$\asymp{\frac{p_{n}}{p_{n+1}}}\prod_{x_{1:n}\in\{0,1\}^{n}:\,i_{x_{1:n}}=1}(N_{x_{1:n}}/{(2\pi)})^{1/2}.$ (C7)
By (30), the term $p_{n}/p_{n+1}$ is strictly positive and bounded, so (C7) gives
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}\to\infty.$ (C8)
Let $\tilde{n}\in\mathcal{N}$ with $\tilde{n}>n^{*}$. By (30), $\{\tilde{n},\tilde{n}-1,\ldots,n^{*}\}\subseteq\mathcal{N}$. Hence, by (C8),
$\frac{p(n^{*}\,|\,Z_{1:m})}{p(\tilde{n}\,|\,Z_{1:m})}=\frac{p(n^{*}\,|\,Z_{1:m})}{p(n^{*}+1\,|\,Z_{1:m})}\cdots\frac{p(\tilde{n}-1\,|\,Z_{1:m})}{p(\tilde{n}\,|\,Z_{1:m})}\to\infty,$
whereby (C2) is verified.
Consider instead any $n\in\mathcal{N}$ with $n<n^{*}$. Define $d_{x_{1:n}}=0$ if $f^{*}$ is constant over $I_{x_{1:n}}$ and $d_{x_{1:n}}=1$ otherwise. Then (C5) writes
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}\asymp{\frac{p_{n}}{p_{n+1}}}$
×$\times\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ d_{x_{1:n}}=0,\,i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}$
×$\times\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ d_{x_{1:n}}=1,\,i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}.$
In a similar way as (C7) was obtained, we see that each term in the first product above is $O((N_{x_{1:n}}/{(2\pi)})^{1/2})$. Consider any $x_{1:n}$ with $d_{x_{1:n}}=1$ and $i_{x_{1:n}}=1$. Defining
$\epsilon_{x_{1:n}}={\int_{I_{x_{1:n},0}}f^{*}(v)\,\mathrm{d}v}\bigg{/}{\int_{I_{x_{1:n}}}f^{*}(v)\,\mathrm{d}v},$
then $\epsilon_{x_{1:n}}\in(0,1)$. By (C3),
$\frac{N_{x_{1:n},0}}{N_{x_{1:n}}}\rightarrow\epsilon_{x_{1:n},0},\quad\frac{N_{x_{1:n},1}}{N_{x_{1:n}}}\rightarrow\epsilon_{x_{1:n},1}=1-\epsilon_{x_{1:n},0}.$
Therefore,
$\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}$
$\qquad=\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{N_{x_{1:n}0}^{N_{x_{1:n},0}-1/2}}{N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}}\frac{N_{x_{1:n}}^{N_{x_{1:n},1}-1/2}}{N_{x_{1:n},1}^{N_{x_{1:n},0}-1/2}}N_{x_{1:n}}/\sqrt{2\pi}$
$\qquad\asymp\left(\frac{\ell_{x_{1:n},0}}{\ell_{x_{1:n}}\epsilon_{x_{1:n},0}}\right)^{N_{x_{1:n}}\epsilon_{x_{1:n},0}}\left(\frac{\ell_{x_{1:n},1}}{\ell_{x_{1:n}}(1-\epsilon_{x_{1:n},0})}\right)^{N_{x_{1:n}}(1-\epsilon_{x_{1:n},0})}$
×$\qquad\qquad\times\epsilon_{x_{1:n},0}^{1/2}(1-\epsilon_{x_{1:n},0})^{1/2}N_{x_{1:n}}^{1/2}/\sqrt{2\pi}$
which is $O(m^{1/2})$ if $\ell_{x_{1:n},0}/\ell_{x_{1:n}}=\epsilon_{x_{1:n},0}$ and converges exponentially fast to 0 otherwise. To prove (C2), since for $\tilde{n}\in\mathcal{N}$ with $\tilde{n}<n^{*}$,
$\frac{p(n^{*}\,|\,Z_{1:m})}{p(\tilde{n}\,|\,Z_{1:m})}=\frac{p(n^{*}\,|\,Z_{1:m})}{p(n^{*}-1\,|\,Z_{1:m})}\cdots\frac{p(\tilde{n}+1\,|\,Z_{1:m})}{p(\tilde{n}\,|\,Z_{1:m})},$
it is sufficient that there exists one $x_{1:j}\in\{0,1\}^{j}$ with $\tilde{n}\leq j\leq n^{*}$ and $\ell_{x_{1:j},0}/\ell_{x_{1:j}}\neq\epsilon_{x_{1:j},0}$. By contradiction, suppose this is false whenever $\tilde{n}\leq j\leq n^{*}$, that is, $\ell_{x_{1:j},0}/\ell_{x_{1:j}}=\epsilon_{x_{1:j},0}$ whenever $x_{1:j}\in\{0,1\}^{j}$ and $\tilde{n}\leq j\leq n^{*}$. Then, for any integer $n\geq\tilde{n}$ and any $x_{1:n}\in\{0,1\}^{n}$,
$\mathrm{P}^{*}(Z_{i}\in I_{x_{1:n}}$ $\,|\,Z_{i}\in I_{x_{1:\tilde{n}}})$
$=\mathrm{P}^{*}(Z_{i}\in I_{x_{1:n}}\,|\,Z_{i}\in I_{x_{1:(n-1)}})\cdots\mathrm{P}^{*}(Z_{i}\in I_{x_{1:(\tilde{n}+1)}}\,|\,Z_{i}\in I_{x_{1:\tilde{n}}})$
$=\epsilon_{x_{1:n}}\epsilon_{x_{1:(n-1)}}\cdots\epsilon_{x_{1:(\tilde{n}+1)}}=\frac{\ell_{x_{1:n}}}{\ell_{x_{1:\tilde{n}}}}.$
Hence, by the $\pi$-$\lambda$ theorem, under the true model $Z_{i}\,|\,Z_{i}\in I_{x_{1:\tilde{n}}}$ is uniformly distributed on $I_{x_{1:\tilde{n}}}$. This is a contradiction since it means that $\tilde{n}$ digits were sufficient under the true model but we have assumed that $\tilde{n}<n^{*}$.
Proof of Theorem 5.5.
Arguing as in the proof of Theorem 5.4 (case $n<n^{*}$), the condition implies that for any $n\in\mathbb{N}_{0}$, almost surely under the true model, $p(n\,|\,Z_{1:m})/p(n+1\,|\,Z_{1:m})\rightarrow 0$ as $m\rightarrow\infty$. Thus the result follows by induction. ∎
Proof of Theorem 5.6.
Define $\tilde{N}_{x_{1:j}}$ with respect to $\tilde{Z}_{1:m}$ in the same way as we defined $N_{x_{1:j}}$ with respect to $Z_{1:m}$ in the proof of Theorem 5.4. Consider any integer $n>n^{*}$. By (20),
$\frac{p(n^{*}\,|\,\tilde{Z}_{1:m})}{p(n\,|\,\tilde{Z}_{1:m})}$ $=\frac{p_{n^{*}}}{p_{n}}\frac{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{\tilde{N}_{x_{1:n}}}}{\prod_{x_{1:n^{*}}\in\{0,1\}^{n^{*}}}\ell_{x_{1:n^{*}}}^{\tilde{N}_{x_{1:n^{*}}}}}$
×$\hskip 56.9055pt\times\frac{1}{\prod_{j=n^{*}+1}^{n}\prod_{x_{1:j}\in\{0,1\}^{j}}\mathrm{b}(\alpha_{x_{1:j},0}+\tilde{N}_{x_{1:j},0},\alpha_{x_{1:j},1}+\tilde{N}_{x_{1:j},1})}.$
By (34), for $j=\bar{n},\bar{n}+1,\ldots$ and every $x_{1:j}\in\{0,1\}^{j}$, we have almost surely under the true distribution that
$\tilde{N}_{x_{1:j}}=\begin{cases}\tilde{N}_{x_{1:\bar{n}}}&\text{if }x_{1:j}=(x_{1:\bar{n}},0,\ldots,0),\\ 0&\text{otherwise}.\end{cases}$
Further, define $A_{j}=\{x_{1:j}\in\{0,1\}^{j}:x_{1:j}=(x_{1:\bar{n}},0,\ldots,0)\text{ for some }x_{1:\bar{n}}\}$. Then,
$\frac{p(n^{*}\,|\,\tilde{Z}_{1:m})}{p(n\,|\,\tilde{Z}_{1:m})}$ $=\frac{p_{n^{*}}}{p_{n}}\prod_{x_{1:\bar{n}}\in\{0,1\}^{\bar{n}}}\left(\frac{\ell_{x_{1:\bar{n}},0,\ldots,0}}{\ell_{x_{1:n^{*}}}}\right)^{\tilde{N}_{x_{1:\bar{n}}}}$
×$\hskip 56.9055pt\times\frac{1}{\prod_{j=\bar{n}+1}^{n}\prod_{x_{1:j}\in A_{j}}\mathrm{b}(\alpha_{x_{1:j},0}+\tilde{N}_{x_{1:j},0},\alpha_{x_{1:j},1})}$
×$\hskip 56.9055pt\times\frac{1}{\prod_{j=\bar{n}+1}^{n}\prod_{x_{1:j}\in\{0,1\}^{j}\setminus A_{j}}\mathrm{b}(\alpha_{x_{1:j},0},\alpha_{x_{1:j},1})}.$
Stirling’s approximation of the beta function gives $\mathrm{b}(x,y)=\Gamma(y)x^{-y}(1+O(1/x))$ for fixed $y>0$. Therefore,
$\frac{p(n^{*}\,|\,\tilde{Z}_{1:m})}{p(n\,|\,\tilde{Z}_{1:m})}=\left[\prod_{x_{1:\bar{n}}\in\{0,1\}^{\bar{n}}}\left(\frac{\ell_{x_{1:\bar{n}},0,\ldots,0}}{\ell_{x_{1:\bar{n}}}}\right)^{\tilde{N}_{x_{1:\bar{n}}}}(N_{x_{1:\bar{n}},0})^{\alpha_{x_{1:j},1}}\right]O(1)$
where the term []delimited-[][$\cdots$][ ⋯ ] goes to 00 as $m\rightarrow\infty$, since for some $x_{1:\bar{n}}$, $\tilde{N}_{x_{1:\bar{n}}}\rightarrow\infty$ as $m\rightarrow\infty$. ∎
Appendix D Proofs for Section 5.3
In the following, for two PDFs $f_{1}$ and $f_{2}$ on $[0,1)$, let
$\mathrm{KL}(f_{1},f_{2})=\int_{0}^{1}f_{1}(x)\log\frac{f_{1}(x)}{f_{2}(x)}\mathrm{d}x$
denote the Kullback-Leibler divergence of $f_{1}$ from $f_{2}$.
Definition D.1 (Kullback-Leibler support).
Let $\Pi$ be a distribution on the set of PDFs on $[0,1)$. We say that a PDF $f$ belongs to the Kullback-Leibler support of $\Pi$ if, for any $\varepsilon>0$,
$\Pi\left(\{g:\mathrm{KL}(f,g)<\varepsilon\}\right)>0.$
Proof of Theorem 5.8.
To prove the statement it is sufficient to show that $f^{*}$ belongs to the Kullback-Leibler support of the prior. This follows from Theorem 6.16 and Example 6.20 in [14].
Consider the case of the GFPT1 prior $\Pi_{1}$, cf. (7). Clearly,
$\Pi_{1}\left(\{f:\mathrm{KL}(f^{*},f)<\varepsilon\}\right)$ $=\sum_{n=0}^{\infty}\Pi_{1}\left(\{f:\mathrm{KL}(f^{*},f)<\varepsilon\}\,|\,N=n\right)p_{n}$
$\geq\Pi_{1}\left(\{f:\mathrm{KL}(f^{*},f)<\varepsilon\}\,|\,N=\bar{n}\right)p_{\bar{n}}$
for any $\bar{n}\in\mathbb{N}_{0}$. Following closely the proof of Theorem 7.1 in [14], we will show that
$\Pi_{1}\left(\{f:\mathrm{KL}(f^{*},f)<\varepsilon\}\,|\,N=\bar{n}\right)>0$
for sufficiently large $\bar{n}$. Setting $y_{x_{1:j-1},k}=\mathrm{P}^{*}(I_{x_{1:j-1},k})/\mathrm{P}^{*}(I_{x_{1:j-1}})$, define the discretization of $f^{*}$ at level $n$ of the NBP $I$ as
$f^{*}_{n}(x)=\prod_{j=1}^{n}y_{x_{1:j-1},0}^{1-x_{j}}y_{x_{1:j-1},1}^{x_{j}}{\ell_{x_{1:j-1}}}/{\ell_{x_{1:j}}}\quad\text{if }x=.x_{1}x_{2}\ldots\in[0,1).$ (D9)
Hence, for almost any $x\in[0,1)$,
$f^{*}(x)=\lim_{n\rightarrow\infty}f^{*}_{n}(x)=\prod_{j=1}^{\infty}y_{x_{1:j-1},0}^{1-x_{j}}y_{x_{1:j-1},1}^{x_{j}}{\ell_{x_{1:j-1}}}/{\ell_{x_{1:j}}}.$
In a similar fashion, rewrite the random PDF in (6) as
$f(x\,|\,Y^{(N)})=\prod_{j=1}^{N}Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}{\ell_{x_{1:j-1}}}/{\ell_{x_{1:j}}}\quad\text{if }x=.x_{1}x_{2}\ldots\in[0,1).$
Then, when $N=\bar{n}$,
$\mathrm{KL}(f^{*},f(\cdot\,|\,Y^{(\bar{n})}))=\int_{0}^{1}f^{*}(x)\log\frac{f^{*}(x)}{f(x\,|\,Y^{(\bar{n})})}\,\mathrm{d}x$
$=\int_{0}^{1}f^{*}(x)\left(\sum_{j=1}^{\bar{n}}\log\frac{y_{x_{1:j-1},0}^{1-x_{j}}y_{x_{1:j-1},1}^{x_{j}}}{Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}}+\sum_{j=\bar{n}+1}^{\infty}\log\left(y_{x_{1:j-1},0}^{1-x_{j}}y_{x_{1:j-1},1}^{x_{j}}\frac{\ell_{x_{1:j-1}}}{\ell_{x_{1:j}}}\right)\right)\mathrm{d}x$
Again from Theorem 7.1 in [14] we see that for any $\delta>0$,
$\Pi_{1}\left(\int_{0}^{1}f^{*}(x)\left(\sum_{j=1}^{\bar{n}}\log\frac{y_{x_{1:j-1},0}^{1-x_{j}}y_{x_{1:j-1},1}^{x_{j}}}{Y_{x_{1:j-1},0}^{1-x_{j}}Y_{x_{1:j-1},1}^{x_{j}}}\right)\mathrm{d}x<\delta\,\bigg{|}\,N=\bar{n}\right)>0.$
On the other hand, if $\bar{n}$ is large enough the term
$\int_{0}^{1}f^{*}(x)\left(\sum_{j=\bar{n}+1}^{\infty}\log\left(y_{x_{1:j-1},0}^{1-x_{j}}y_{x_{1:j-1},1}^{x_{j}}\frac{\ell_{x_{1:j-1}}}{\ell_{x_{1:j}}}\right)\right)\mathrm{d}x$
is negligible by the assumption $\int f^{*}(x)\log f^{*}(x)\,\mathrm{d}x<\infty$ and Lemma B.10 in [14]. Hence, $f^{*}$ belongs to the Kullback-Leibler support of $\Pi_{1}$, whereby the proof is complete in case of the GFPT1 prior.
The proof under the GFPT2 prior follows by observing that conditioned on $R$, the GFPT2 prior coincides with a GFPT1 prior, and then applying Proposition 6.28 in [14]. ∎
For the proof of Theorem 5.9, we need some additional notation. For a (semi)metric space $(S,d)$ and a totally bounded subset $C\subset S$, define the $\varepsilon$-covering number of $C$ as the smallest integer of balls of radius $\varepsilon$ (with respect to $d$) needed to cover $C$. Denote this $\varepsilon$-covering number by $N(\varepsilon,C,d)$. The proof of Theorem 5.9 is based on the following general result, where $\mathcal{P}$ denotes the set of absolutely continuous probability distributions on $[0,1)$ and this set is equipped with the Borel $\sigma$-algebra generated by the Hellinger distance.
Theorem D.2 (Theorem 6.23 in [14]).
Let $\Pi\in\mathcal{P}$. Consider the data model $Z_{1},\ldots,Z_{m}\,|\,g\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}g$ with prior PDF $g\sim\Pi$, and denote by $\Pi(\cdot\,|\,Z_{1:m})$ the associated posterior distribution. Assume that the data were generated i.i.d. from a probability measure $\mathrm{P}^{*}$ with a PDF $f^{*}$ which belongs to the Kullback-Leibler support of $\Pi$, and for every $\delta>0$, there exist partitions $\mathcal{P}=\mathcal{P}_{m,1}\cup\mathcal{P}_{m,2}$ and a constant $C>0$ such that for all sufficiently large values of $m$,
  • $(i)$
    $\log N(\delta,\mathcal{P}_{m,1},d_{\mathrm{H}})\leq m\delta^{2}$,
  • $(ii)$
    $\Pi(\mathcal{P}_{m,2})\leq\mathrm{e}^{-Cm}$.
Then, for any $\varepsilon>0$,
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}\Pi(\{f\colon d_{\mathrm{H}}(f,f^{*})<\varepsilon\}\,|\,Z_{1:m})=1\right)=1.$
Proof of Theorem 5.9.
We apply Theorem D.2. Since $f^{*}$ belongs to the Kullback-Leibler support of the prior, cf. the proof of Theorem 5.8, we just need to verify items $(i)$ and $(ii)$ in Theorem D.2. To do so, we follow closely the proof of Theorem 7.16 in [14]: For a given PDF $f$ on $[0,1)$, denote by $f_{k}$ its discretization at level $k$ of the NBP $I$, cf. (D9). Consider partitions
$\mathcal{P}_{m,1}=\{f\colon\|\log(f/f_{k_{m}})\|_{\infty}\leq\varepsilon^{2}/8\},\quad\mathcal{P}_{m,2}=\mathcal{P}\setminus\mathcal{P}_{m,1},$
such that $A\log m\leq k_{m}\leq B\log m$ where $0<A<B$ are constants. Then the proof of item $(i)$ above follows the same lines as the proof of Theorem 7.16 in [14]. Concerning item $(ii)$, write
$\Pi_{1}\left(f(\cdot\,|\,Y^{(N)})\in\mathcal{P}_{m,2}\right)=\sum_{n=0}^{\infty}p_{n}\Pi_{1}\left(f(\cdot\,|\,Y^{(N)})\in\mathcal{P}_{m,2}\,\big{|}\,N=n\right)$
and observe that $\Pi_{1}(\cdot\,|\,N=n)$ gives positive probability only to those densities that are piecewise constant after level $n$ of the NBP $I$. Thus,
$\Pi_{1}\left(f(\cdot\,|\,Y^{(N)})\in\mathcal{P}_{m,2}\right)=\sum_{n=k_{m}+1}^{\infty}p_{n}\Pi_{1}\left(f(\cdot\,|\,Y^{(N)})\in\mathcal{P}_{m,2}\,\big{|}\,N=n\right)\leq De^{-ck_{m}}$
for some constant $D$ depending only on $C$ and $c$, cf. Theorem 5.9. Hence, item (ii) above follows by taking $\delta=\delta(m)=m^{1/2}\log m$. ∎